### v2

# Under-coverage in high-statistics counting experiments with finite MC samples

Abstract
We consider the problem of setting confidence intervals on a parameter of interest from the maximum-likelihood fit of a physics model to a binned data set with a large number of bins, large event-counts per bin, and in the presence of systematic uncertainties modeled as nuisance parameters. We use the profile-likelihood ratio for statistical inference and focus on the case in which the model is determined from Monte Carlo simulated samples of finite size. We start by presenting a toy model in which the asymptotic properties of the profile-likelihood ratio are manifestly broken even if the numbers of events per bin in both the data and simulated samples are seemingly large enough to warrant their validity. We then move to the general setting to show how statistical uncertainties in the Monte Carlo predictions can affect the coverage of confidence intervals constructed in the asymptotic approximation always in the same direction, namely they lead to systematic under-coverage.
arxiv: 2401.10542
1 Introduction
Experiments in particle physics are often concerned with the problem of estimating a parameter of interest ($\mu$) from the maximum-likelihood fit of a physics model to sampling distributions constructed from a large number of identically repeated events, such as decays or collisions of particles. When the number of events is large enough to populate the full space of experimental outcomes, sampling distributions can be more conveniently approximated by histograms with a finite number of bins, resulting in a so-called binned analysis. Best-fit values for the parameter of interest (POI) are generally of limited use unless presented together with a confidence interval (CI), that is the range of allowed values in which the true value $\mu_{\rm t}$ of the POI is likely to be found within some conventional level of confidence.
Let us then assume that the experiment under consideration aims at measuring the number of occurrences of some particle physics reaction in the $n$ bins of a histogram, and that the expected number of events in each bin is related to $\mu$ according to a given physics model. Under fairly general assumptions, the event counts ${\bf y}=(y_{1},\ldots,y_{n})$ can be modeled as a vector of $n$ independent and Poisson-distributed random variables whose expectation value $\langle{\bf y}\rangle={\bf f}(\mu)$ is thus sufficient to fully define the joint probability density function of the data. The model ${\bf f}(\mu)$ might be affected by other sources of systematic uncertainties, related for example to imperfect knowledge of the detector response, the presence of backgrounds, or additional unknown parameters. We will assume that these uncertainties can be accounted for by extending the parameter space to include additional variables $\boldsymbol{\theta}$, commonly known as nuisance parameters.
Due to the complexity of modeling particle physics processes from generator up to detector-level, the function ${\bf f}(\mu,\boldsymbol{\theta})$ can rarely be computed analytically, leaving Monte Carlo (MC) techniques as the only viable way for its determination. Given that MC samples necessarily comprise a finite number of events, a MC-simulated model prediction will never be exact, thus introducing an additional source of systematic uncertainty. A general rule is that the size of the samples should exceed that of the data by at least a few times. However, such a request can turn out computationally expensive or even unfeasible for analyses performed on high-statistics samples of events, especially when CPU-intensive MC simulations are needed to achieve the desired level of theoretical accuracy. In such cases, one has to live with samples of MC-simulated events of size comparable or smaller than the data. The situation can become even more problematic when the simulated events are weighted Glusenkamp .
The correct frequentist approach to deal with the finite size of MC samples was first introduced by Barlow and Beeston in ref. BARLOW1993219 . In their seminal paper, the authors proposed the idea of treating the uncertain MC predictions as additional nuisance parameters of which the available MC sample represents an ancillary measurement. Since then, several improvements in the treatment of MC uncertainties have been proposed in the literature, with particular emphasis on the case of low-statistic samples and the proper handling of event weights BARLOW1993219 ; chirkin2013likelihooddescriptioncomparingdata ; Glusenkamp ; Arguelles .
In this paper, we focus on the problem of setting confidence intervals on a parameter of interest in the presence of nuisance parameters and in the case of high-statistics counting experiments, for which asymptotic formulas from the theory of maximum likelihood estimators Cowan ; Kendall ; Asimov are expected to hold. We notice that the high-statistics case has received less attention in the literature compared to the low-statistics limit, where asymptotic formulas are more likely invalidated by low event counts. Furthermore, the high-statistics case is presently of great interest for precision measurements performed on large-scale data samples, such as those collected at the LHC, see e.g. refs. ATLAS:2024wla ; CMS:2024lrd for a recent review on this subject.
A relevant example of a precision measurement performed on a large-scale data sample, and with a large number of nuisance parameters, is represented by the measurement of the $W$ boson mass by the CMS Collaboration CMS:2024lrd : the CMS measurement analyzed a sample of LHC collision data consisting of about $10^{8}$ single-muon events, binned in the kinematics and charge of the muons, resulting in approximately $3\times 10^{3}$ bins, which are used to simultaneously measure the mass of the $W$ boson together with a few thousands of nuisance parameters via a maximum-likelihood fit. The model is constructed by using samples of MC-simulated events whose overall statistical power exceeds that of the data by a factor of about $4$. Alternate MC predictions corresponding to ±$\pm 1\sigma$ variations of each nuisance parameter are constructed by reweighting the nominal MC simulation. The effect of statistical fluctuations in the alternate templates used to construct the systematic variations was observed to lead to a moderate under-coverage of Hessian uncertainties, which required an inflation of the statistical uncertainty of the MC templates to restore the correct coverage CMS:2024lrd . The main scope of this paper is to study the origin of such an effect from a more general perspective.
This paper is organized as follows. We start by presenting a hypothetical experiment which can be considered paradigmatic of our case of interest: the data set consists in a histogram with a large number of bins; the event counts in each bin are large; the physics model used to describe the data includes nuisance parameters and is determined from MC-simulated samples of size comparable to the data. We compute numerically the coverage of intervals constructed from CI-setting methods that rely on the asymptotic properties of maximum-likelihood estimators, such as those based on the Hessian matrix or the profile-likelihood ratio, and find that such intervals cover systematically less than their claimed confidence level. Then, we try alternative CI-setting methods that do not rely on asymptotic properties, such as those inspired by the unified approach by Feldman and Cousins FC and again find similar issues. Inspired by our findings, we propose a heuristic definition of a confidence interval that better approaches the correct Feldman-Cousins interval. In the second part of the paper, we study the same problem from a more general perspective to ascertain in which circumstances statistical fluctuations of the MC model can lead to a systematic under-coverage of asymptotic confidence intervals. We conclude by suggesting numerical tests to quantify the impact of MC statistical fluctuations for a generic problem, remarking that the assumption of being in the asymptotic regime may not be simply guaranteed by the large size of the data and simulated samples alone.
2 A paradigmatic toy model
Consider a hypothetical experiment where two processes, respectively labeled with index $j=1,2$, generate independent, Poisson-distributed event counts $y_{ji}$ with expectation values in the $i^{\rm th}$ bin given by
$\langle y_{1i}\rangle\equiv\nu_{1i}=K\;\;\;{\rm and}\;\;\;\langle y_{2i}\rangle\equiv\nu_{2i}=\begin{cases}K\cdot(1+\epsilon)&\;\;{\rm for}\;i<\frac{n}{2}\\ K&\;\;{\rm for}\;i\geq\frac{n}{2}\end{cases}$ (1)
where $|\epsilon|\ll 1$, $K\gg 1$, and $n$ is the (even) number of analyzed bins. In a real experiment the constant $K$ would be related to e.g. beam flux, detector efficiency, cross sections, branching ratios, etc. In our study, we find it more convenient to trade it for the total number of expected events $N$ in the data sample.
Let the purpose of this hypothetical experiment be a measurement of the total number of events from the process $j=1$, which can be thought of as the “signal”, with process $j=2$ playing the role of a “background” of unknown cross section. It is easy to see how the data could be reduced to just two bins with no loss of information on the parameter of interest. However, the crucial assumption made here is that the true model of eq. (1) is not known a priori but can only be predicted from finite-size MC samples consisting of ${t}_{ji}$ simulated events in the $i^{\rm th}$ bin and for the $j^{\rm th}$ process. Since $t_{ji}$ are also random variables, their values will differ from $\nu_{ij}$ by statistical fluctuations introduced by the MC generation, which could indeed justify preserving the full granularity of the data.
Given a set of MC predictions $t_{ji}$, the signal normalization can be extracted by parameterizing the expectation value $\langle y_{i}\rangle$ as:
$f_{i}(\mu_{1},\mu_{2}\,;\,t_{1i},t_{2i})=\sum_{j=1}^{2}\mu_{j}\frac{t_{ji}}{k_{j}}$ (2)
where strength modifiers $\mu_{j}$ have been introduced for the two processes as a proxy for the total number of expected events, and $k_{j}$ are constants that rescale the normalization of the MC samples to some reference value. Rescaled MC predictions $T_{ji}=t_{ji}/k_{j}$, commonly known as templates, provide the elements of the Jacobian matrix of ${\bf f}$ with respect to its parameters. We choose $k_{j}$ equal to the ratio between the expectation value of $t_{ji}$ and the true value $\nu_{ji}$, so that by construction $f_{i}$ is an unbiased estimator of $\langle y_{i}\rangle$ for $\mu_{1}\equiv\mu_{1,{\rm t}}=1$ and $\mu_{2}\equiv\mu_{{2,\rm t}}=1$. Consistently with this assumption, values for $y_{ji}$ and $t_{ji}$ can be generated by drawing random integer numbers from the Poisson density with mean values $\nu_{ji}$ and $k_{j}\nu_{ji}$, respectively, that is:
${\rm Prob}\left[y_{ij}\right]=\frac{e^{-\nu_{ji}}\left(\nu_{ji}\right)^{y_{ji}}}{y_{ij}!}\;\;\;{\rm and}\;\;\;{\rm Prob}\left[t_{ij}\right]=\frac{e^{-k_{j}\nu_{ji}}\left(k_{j}\nu_{ji}\right)^{t_{ji}}}{t_{ij}!}$ (3)
Examples of templates $T_{ji}$ are shown in figure 1 for representative values of $N$, $\epsilon$, $n$, and $k_{j}$.
Figure 1: Examples of templates $T_{ji}$ generated as described in the text for $N=2\times 10^{6}$, $\epsilon=0.03$, $n=200$, and for two values of $k_{1}=k_{2}=k$, namely $k=1$ (left) and $k=10$ (right).
[…]
The coverage of $I$ is defined as the fraction of pseudo-experiments for which $\mu_{\rm t}\in I$. In order to verify that the coverage of $I$ does not depend on $\mu_{\rm t}$, two additional independent sets of data ${\bf y}_{\pm}$ are considered. The latter are generated as in eq. (3) with the replacement
$\nu_{1i}\to\nu^{\pm}_{1i}=\nu_{1i}\cdot\left(1\pm 5\bar{\sigma}_{\rm H}\right),$ (4)
where $\bar{\sigma}_{{\rm H}}$ is defined as the standard deviation of $\hat{\mu}$ when the MC templates $T_{ji}$ are replaced by their true values $\nu_{ji}$, or, equivalently, in the limit of infinite-size MC samples, that is $k_{j}\to\infty$. Clearly, $\mu_{\rm t}=1\pm 5\bar{\sigma}_{{\rm H}}$ for these alternate data sets.
2.1 Likelihood functions for the toy model
The total event counts per bin $y_{i}$ can be treated as $n$ independent random variables drawn from the Poisson density $\mathcal{P}$, leading to the definition of the likelihood as:
$L_{\rm P}(\mu,\theta)=\prod_{i=1}^{n}\mathcal{P}(y_{i}\,|\,f_{i}(\mu,\theta)).$ (5)
The maximum-likelihood estimator (MLE) for the likelihood parameters is the point $(\hat{\mu},\hat{{\theta}})$ that minimizes the negative log-likelihood function (NLL), which, apart from an irrelevant constant, can be written as
$-\ln L_{\rm P}(\mu,\theta)=\sum_{i=1}^{n}\left[f_{i}(\mu,\theta\,;\,t_{1i},t_{2i})-y_{i}\ln f_{i}(\mu,\theta\,;\,t_{1i},t_{2i})\right].$ (6)
Equation (6) neglects the fact that $t_{ji}$ are also random variables. A straightforward way to account for their stochastic nature is the Barlow-Beeston approach BARLOW1993219 , which consists in treating the vectors ${\bf t}_{j}=(t_{j1},\ldots,t_{jn})$ as independent Poisson-distributed random variables with unknown mean values $\boldsymbol{\theta}_{j}$, therefore extending the likelihood to include an extra contribution:
$-\ln L_{\rm P+BB}(\mu,\theta,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$ $=\sum_{i=1}^{n}\left[f_{i}(\mu,\theta\,;\,\theta_{1i},\theta_{2i})-y_{i}\ln f_{i}(\mu,\theta\,;\,\theta_{1i},\theta_{2i})\right]$
$+\sum_{i=1}^{n}\sum_{j=1}^{2}\left(\theta_{ji}-t_{ji}\ln\theta_{ji}\right).$ (7)
A major drawback of eq. (7) is the growth in dimensionality: the number of likelihood parameters increases from $p=2$ to $p=2(n+1)$. However, as discussed in ref. BARLOW1993219 , numerical minimization of eq. (7) can be still recast as a two-dimensional optimization problem, at the cost of solving a non-linear system of equations BARLOW1993219 at each step of the minimization, which can be challenging numerically BBlite . For this reason, a simplification to eq. (7) has been proposed by Cowan BBlite , leading to the so-called Barlow-Beeston lite likelihood:
$-\ln L_{\rm P+BBlite}(\mu,\theta,\boldsymbol{\beta})=$
$\sum_{i=1}^{n}\left[f_{i}(\mu,\theta\,;\,\beta_{i}t_{1i},\beta_{i}t_{2i})-y_{i}\ln f_{i}(\mu,\theta\,;\,\beta_{i}t_{1i},\beta_{i}t_{2i})\right]$
$+\sum_{i=1}^{n}\beta_{i}m_{i}-m_{i}\ln\beta_{i},$ (8)
where the last term accounts for the total statistical uncertainty of the MC simulation in the $i^{\rm th}$ bin as represented by the effective number of MC events $m_{i}$ defined as
$m_{i}=\left(\frac{t_{1i}}{k_{1}}+\frac{t_{2i}}{k_{2}}\right)^{2}\left(\frac{t_{1i}}{k_{1}^{2}}+\frac{t_{2i}}{k_{2}^{2}}\right)^{-1}.$ (9)
Equation (2.1) still implies an increase in dimensionality, but this time the additional parameters $\boldsymbol{\beta}$ can be profiled analytically with a negligible computational overhead. This feature, together with its capability of catching the dominant effect of MC uncertainties, has made the Barlow-Beeston lite approach a very popular and successful solution in HEP analyses.
2.1.1 Alternate likelihood parametrization
A data set distributed according to eq. (1) is clearly not sensitive to the signal normalization when $\epsilon=0$, which translates into a degeneracy of the likelihood function with respect to its parameters. For $\epsilon>0$, the problem is mathematically well defined, but the separation between the two processes becomes increasingly more difficult as $\epsilon$ approaches zero due to the large anti-correlation between $\mu_{1}$ and $\mu_{2}$. For example, for values of the $\epsilon$ parameter of ${\cal O}(1\%)$, the linear correlation $\rho$ between the two parameters is almost complete, e.g. $\rho=-0.99989$ for $\epsilon=0.03$. In fact, one might argue that the individual measurement of $\mu_{1}$ and $\mu_{2}$ is perhaps a poorly defined problem, given that the data is much more sensitive to their sum than to the difference. The role played by the choice of parameters is investigated by considering an alternate parametrization of the likelihood defined by the linear transformation
$\mu_{1,2}^{\prime}=\frac{1}{2}\left(\mu_{1}\mp\mu_{2}\right),$ (10)
which rotates the basis of the original parameters closer to the eigenvectors of the Hessian matrix. In fact, after applying the transformation in eq. (10), the linear correlation coefficient [^2] We remark that the condition number of the covariance matrix is invariant under a linear transformation such as eq. (10), which indeed offers a better metric to characterize the problem. between the transformed parameters becomes $0.71$ for the same value of $\epsilon$. However, we remark that the reduced correlation between parameters is not for free, given that the effective MC statistical power of the transformed templates $T_{ji}^{\prime}$ has also changed. The obvious solution would be to directly generate templates $T_{ji}^{\prime}$ for the transformed parameters, although this might not always be feasible for generic problems, for example because the MC generator is not flexible enough to allow for it.
When considering the alternate parametrization of eq. (10), we choose the POI (NP) to be $\mu_{1}^{\prime}$ ($\mu_{2}^{\prime}$) and change the definition of the alternate models accordingly, that is:
$\nu^{\pm}_{1i}=\nu_{1i}\cdot\left(1\pm 5\bar{\sigma}_{{\rm H}}^{\prime}\right),\;\;\;\nu^{\pm}_{2i}=\nu_{2i}\cdot\left(1\mp 5\bar{\sigma}_{{\rm H}}^{\prime}\right),$ (11)
where $\bar{\sigma}_{{\rm H}}^{\prime}$ is defined as the standard deviation of $\hat{\mu}_{1}^{\prime}$ for the nominal model in the limit of infinite-size MC samples. Given that $\mu_{1}$ and $\mu_{2}$ are highly correlated, it is easy to see that $\bar{\sigma}_{{\rm H}}^{\prime}\approx\bar{\sigma}_{{\rm H}}$, so for the sake of simplicity we will set $\bar{\sigma}_{{\rm H}}^{\prime}=\bar{\sigma}_{{\rm H}}$ in eq. (11). For this other parametrization, we have $\mu_{\rm t}=0$ for the nominal data and $\mu_{\rm t}=\pm 5\bar{\sigma}_{\rm H}$ for the alternate.
The experiment can now be interpreted as measuring simultaneously the total number of events together with the asymmetry ${\cal A}$ between the two halves of the histogram. Indeed, by denoting the number of expected events in the first (second) half of the bins as $N_{<}$ ($N_{>}$), it is easy to prove that for the true model
${\cal A}=\frac{N_{<}-N_{>}}{N_{<}+N_{>}}=\frac{\epsilon}{4}\left(1-\frac{\mu_{1}^{\prime}}{\mu_{2}^{\prime}}+{\cal O}(\epsilon)\right).$ (12)
The statistical uncertainty on $\mu_{2}^{\prime}$ is of ${\cal O}(N^{-1/2})$, so it will be highly constrained when $N$ is large. Hence, the uncertainty on ${\cal A}$, which is directly proportional to the uncertainty on the ratio $\mu^{\prime}_{1}/\mu_{2}^{\prime}$ as for eq. (12), is mostly driven by the limited precision on $\mu_{1}^{\prime}$. Measurements of asymmetries between two portions of a distribution occur frequently in Particle Physics. The determination of the $W$ boson mass at hadron colliders from the distribution of transverse momenta of the decaying leptons, already mentioned in section 1, is a relevant example of this kind. Indeed, this measurement can be interpreted, at least to a first approximation, as the determination of the asymmetry of the transverse momentum distribution around the Jacobian peak Rottoli:2023xdc .
2.1.2 Likelihood functions in the Gaussian approximation
In the limit $y_{i}\gg 1$, numerical approximations of the negative log-likelihood function can be made. Indeed, the latter consists in a sum of terms of the same form which can be expanded around $f_{i}=y_{i}$ yielding:
$f_{i}-y_{i}\ln f_{i}=\frac{\left(f_{i}-y_{i}\right)^{2}}{y_{i}}+\frac{2}{3}\frac{\left(f_{i}-y_{i}\right)^{3}}{y_{i}^{2}}+\frac{1}{2}\frac{\left(f_{i}-y_{i}\right)^{4}}{y_{i}^{3}}+\ldots,$ (13)
where the dots indicate terms that either do not depend on $f_{i}$ or are of higher-order in the expansion. Neglecting terms beyond the quadratic order amounts to approximate the joint Poisson density as a Gaussian multivariate with covariance matrix ${\bf V}={\rm diag}({\bf y})$. For the toy model, it has been verified that a joint Poisson density and its Gaussian approximation result in consistent point estimators and confidence intervals for all values of $N$ and $n$ considered in this study. For example, for $N=2\times 10^{6}$ and $n=200$, the number of events per bin is $y_{i}\sim 10^{4}$, and the second term in the expansion represents a correction of order $(y_{i}-\nu_{i})/y_{i}\sim\mathcal{O}(\sqrt{y_{i}})/y_{i}\sim 1\%$. The complete list of likelihood functions in the Gaussian approximation used in the study is presented in appendix A.
Another practical advantage of the Barlow-Beeston lite approach discussed in section 2.1 becomes evident in the Gaussian approximation. In order to see this, we first notice that the right-hand side of eq. (6) converges to a quadratic function of ${\bf f}$ in the limit $f_{i}\sim y_{i}\gg 1$, with Hessian matrix given by ${\bf V}^{-1}$. Profiling the $\boldsymbol{\beta}$ parameters in eq. (2.1) leads instead to a non-linear function of $(\theta,\mu)$, see e.g. eq. (72) in the appendix. However, the profiled likelihood function is still approximately quadratic in ${\bf f}$, with Hessian matrix given by $({\bf V}+{\bf V}_{\rm MC})^{-1}$, where ${\bf V}_{\rm MC}$ is a diagonal matrix with $i^{\rm th}$ diagonal-element equal to $f^{2}_{i}/m_{i}\approx y_{i}/k$, where $k$ is the relative proportion between data and simulation. This observation allows for an effective way to approximate the Barlow-Beeston lite likelihood as a quadratic function of its parameters, henceforth named “Gauss + MC stat.”, through a simple rescaling ${\bf V}\to{\bf V}\cdot(1+k^{-1})$. Numerical differences between minimizing the Barlow-Beeston lite likelihood or its quadratic approximation are found to be negligible in this context.
2.2 Confidence intervals in the asymptotic limit
Based on the properties of the maximum-likelihood estimators and of the profile-likelihood ratio in the asymptotic limit Kendall , we consider the following methods to set confidence intervals on $\mu$.
  • Hessian uncertainty
    The Hessian uncertainty $\hat{\sigma}_{\rm H}$ is defined as the square root of the appropriate diagonal element of the inverse Hessian matrix of the NLL function at the minimum:
    ${\bf C}=\left(-\frac{\partial^{2}\ln L}{\partial^{2}(\mu,\boldsymbol{\theta})}\right)^{-1},$ (14)
    where $\boldsymbol{\theta}$ is a shorthand for the complete set of nuisance parameters. In the asymptotic limit, the maximum-likelihood estimators are known to be Gaussian-distributed around their true values Cowan , with covariance matrix given by the expectation value of ${\bf C}$, so that
    $\left[\hat{\mu}-\hat{\sigma}_{\rm H},\;\hat{\mu}+\hat{\sigma}_{\rm H}\right],$ (15)
    is (asymptotically) a 68.3% CL interval for $\mu$.
  • Profile-Likelihood ratio
    The profile-likelihood ratio (PLR) is a function of $\mu$ defined as Kendall
    $t_{\mu}=-2\ln\frac{L(\mu,{\hat{\boldsymbol{\theta}}}_{\mu})}{L(\hat{\mu},{\hat{\boldsymbol{\theta}}})},$ (16)
    where ${\hat{\boldsymbol{\theta}}}_{\mu}$ is the NP point that minimizes the NLL at a fixed value of $\mu$. In the asymptotic limit, the distribution of $t_{\mu}$ tends to a chi-square with one degree of freedom as for Wilks’ theorem Wilks , with deviations expected at the ${\cal O}(1/\sqrt{N})$ level, with $N$ related to the size of the data. If this high-order correction can be neglected, the interval
    $\{\mu\,:\,t_{\mu}\leq 1\}.$ (17)
    is again a 68.3% CL interval for $\mu$.
Given that $\hat{\mu}$ is also asymptotically Gaussian-distributed, it follows that the confidence intervals defined by eq. (14) and (16) coincide in the asymptotic limit.
2.3 Numerical results
The minimization of the negative log-likelihood function is performed numerically using the MIGRAD minimizer implemented in Minuit2 JAMES1975343 , with PLR scans performed by MINOS and Hessian-based uncertainties computed by HESSE. Linear algebra operations are performed using the Eigen library eigenweb in double precision. The nominal configuration parameters for the toy model are reported in table 1. The effect of changing their values is studied in section 2.5. For simplicity, we have assumed $k_{1}=k_{2}=k$ [^3] For example, this would be appropriate when the signal and background can be defined as sub-processes of a same MC sample. The generalization to arbitrary $k_{j}$ is also straightforward and is not a critical aspect for this study.. Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu_{1}$ are obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments, and reported in table 2, while the results for the other choice $\mu=\mu^{\prime}_{1}$ are shown in table 3.
The expected Hessian uncertainty $\bar{\sigma}_{\rm H}$ for the nominal model and in the limit of infinite-size MC samples, that is when $T_{ji}=\nu_{ji}$, can be computed analytically. Using well-known results, which will be further elaborated in section 3, it is easy to prove that
$\bar{\sigma}_{\rm H}=\frac{4}{|\epsilon|\sqrt{N}}\left(1+{\cal O}(\epsilon)\right)\approx 0.094,$ (18)
in very good agreement with the numerical result shown in the first row of tables 2 and 3. Also, one can easily verify that the linear correlation between $\mu_{1}$ and $\mu_{2}$ differs from $-1$ by terms of ${\cal O}(\epsilon^{2})$. Differences between the Hessian uncertainty for the nominal and alternate models in table 2 are mostly due to the different number of expected events when $\mu_{\rm t}$ is changed.
$N\,\left[10^{6}\right]$ $\epsilon\,\left[\%\right]$ $n$ $k_{1}=k_{2}=k$
${\bf 2}$, $10$, $100$ ${\bf 3}$, $6$, $12$ $20$, $100$, ${\bf 200}$ ${\bf 1}$, $10$, $40$
Table 1: Nominal (in bold) and alternate configuration parameters for the toy model.
Likelihood Minimim. CI method $\mu_{\rm t}-1$ Coverage $\hat{\sigma}$ (mean, median)
1) Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.095, $,\;0.095$ 0.095
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.106, $,\;0.106$ 0.106
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.083, $,\;0.083$ 0.083
2) Poisson Numeric Hessian 0 $0.521(5)$ 0.058, 0.0570.0580.0570.058,\;0.0570.058 , 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.160(4)$ 0.064, $,\;0.064$ 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.064(2)$ 0.050, $,\;0.050$ 0.050
3) Gauss Numeric Hessian 0 $0.520(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.161(4)$ 0.064, $,\;0.064$ 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.063(2)$ 0.050, $,\;0.050$ 0.050
4) Gauss Analytic Hessian 0 $0.520(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.161(4)$ 0.064, $,\;0.064$ 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.063(2)$ 0.050, $,\;0.050$ 0.050
5) Gauss + MC stat. Analytic Hessian 0 ${\bf 0.682}(5)$ 0.081, $,\;0.081$ 0.081
$+5\bar{\sigma}_{{\rm H}}$ $0.230(4)$ 0.086, 0.0850.0860.0850.086,\;0.0850.086 , 0.085
$-5\bar{\sigma}_{{\rm H}}$ $0.124(3)$ 0.076, $,\;0.076$ 0.076
6) Gauss + BB-lite Numeric Hessian 0 ${\bf 0.679}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ $0.259(4)$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.097(3)$ 0.066, $,\;0.066$ 0.066
7) Gauss + BB Numeric Hessian 0 $0.439(5)$ 0.137, 0.1320.1370.1320.137,\;0.1320.137 , 0.132
$+5\bar{\sigma}_{{\rm H}}$ $0.449(5)$ 0.163, 0.1570.1630.1570.163,\;0.1570.163 , 0.157
$-5\bar{\sigma}_{{\rm H}}$ $0.454(5)$ 0.116, 0.1110.1160.1110.116,\;0.1110.116 , 0.111
8) Gauss + BB Numeric PLR 0 $0.454(5)$ 0.141, 0.1350.1410.1350.141,\;0.1350.141 , 0.135
$+5\bar{\sigma}_{{\rm H}}$ $0.461(5)$ 0.170, 0.1620.1700.1620.170,\;0.1620.170 , 0.162
$-5\bar{\sigma}_{{\rm H}}$ $0.469(5)$ 0.116, 0.1110.1160.1110.116,\;0.1110.116 , 0.111
9) Gauss + BB Numeric sPFC 0 $0.471(5)$ 0.160, 0.1460.1600.1460.160,\;0.1460.160 , 0.146
$+5\bar{\sigma}_{{\rm H}}$ $0.464(5)$ 0.184, 0.1690.1840.1690.184,\;0.1690.184 , 0.169
$-5\bar{\sigma}_{{\rm H}}$ $0.614(5)$ 0.166, 0.1550.1660.1550.166,\;0.1550.166 , 0.155
10) Gauss + BB Numeric PFC 0 $0.484(5)$ 0.172, 0.1570.1720.1570.172,\;0.1570.172 , 0.157
$+5\bar{\sigma}_{{\rm H}}$ $0.546(5)$ 0.211, 0.1880.2110.1880.211,\;0.1880.211 , 0.188
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.656}(5)$ 0.176, 0.1600.1760.1600.176,\;0.1600.176 , 0.160
11) Gauss + BB Numeric Barlett 0 $0.480(5)$ 0.168, 0.1550.1680.1550.168,\;0.1550.168 , 0.155
$+5\bar{\sigma}_{{\rm H}}$ $0.513(5)$ 0.205, 0.1840.2050.1840.205,\;0.1840.205 , 0.184
$-5\bar{\sigma}_{{\rm H}}$ $0.615(5)$ 0.166, 0.1520.1660.1520.166,\;0.1520.166 , 0.152
12) Gauss + BB Numeric CH 0 ${\bf 0.681}(5)$ 0.236, 0.2210.2360.2210.236,\;0.2210.236 , 0.221
$+5\bar{\sigma}_{{\rm H}}$ $0.776(4)$ 0.296, 0.2830.2960.2830.296,\;0.2830.296 , 0.283
$-5\bar{\sigma}_{{\rm H}}$ $0.847(4)$ 0.281, 0.2760.2810.2760.281,\;0.2760.281 , 0.276
13) Gauss + BB Numeric FC Cheat 0 ${\bf 0.664}(5)$ 0.229, 0.2070.2290.2070.229,\;0.2070.229 , 0.207
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.673}(5)$ 0.258, 0.2360.2580.2360.258,\;0.2360.258 , 0.236
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.680}(5)$ 0.190, 0.1710.1900.1710.190,\;0.1710.190 , 0.171
14) Gauss + BB Numeric Heuristic 0 ${\bf 0.678}(5)$ 0.228, 0.2210.2280.2210.228,\;0.2210.228 , 0.221
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.660}(5)$ 0.256, 0.2480.2560.2480.256,\;0.2480.256 , 0.248
$-5\bar{\sigma}_{{\rm H}}$ $0.720(4)$ 0.204, 0.1970.2040.1970.204,\;0.1970.204 , 0.197
Table 2: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
1) Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
2) Poisson Numeric Hessian 0 $0.516(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, $,\;0.057$ 0.057
$-5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, $,\;0.057$ 0.057
3) Gauss Numeric Hessian 0 $0.516(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, $,\;0.057$ 0.057
$-5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, $,\;0.057$ 0.057
4) Gauss Analytic Hessian 0 $0.516(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, $,\;0.057$ 0.057
$-5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, $,\;0.057$ 0.057
5) Gauss + MC stat. Analytic Hessian 0 ${\bf 0.675}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, $,\;0.080$ 0.080
$-5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, $,\;0.080$ 0.080
6) Gauss + BB-lite Numeric Hessian 0 ${\bf 0.674}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, $,\;0.080$ 0.080
$-5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.081, 0.0800.0810.0800.081,\;0.0800.081 , 0.080
7) Gauss + BB Numeric Hessian 0 $0.443(5)$ 0.139, 0.1330.1390.1330.139,\;0.1330.139 , 0.133
$+5\bar{\sigma}_{{\rm H}}$ $0.468(5)$ 0.147, 0.1400.1470.1400.147,\;0.1400.147 , 0.140
$-5\bar{\sigma}_{{\rm H}}$ $0.465(5)$ 0.147, 0.1410.1470.1410.147,\;0.1410.147 , 0.141
8) Gauss + BB Numeric PLR 0 $0.448(5)$ 0.140, 0.1340.1400.1340.140,\;0.1340.140 , 0.134
$+5\bar{\sigma}_{{\rm H}}$ $0.474(5)$ 0.152, 0.1440.1520.1440.152,\;0.1440.152 , 0.144
$-5\bar{\sigma}_{{\rm H}}$ $0.470(5)$ 0.143, 0.1370.1430.1370.143,\;0.1370.143 , 0.137
9) Gauss + BB Numeric sPFC 0 $0.474(5)$ 0.158, 0.1440.1580.1440.158,\;0.1440.158 , 0.144
$+5\bar{\sigma}_{{\rm H}}$ $0.597(5)$ 0.211, 0.2000.2110.2000.211,\;0.2000.211 , 0.200
$-5\bar{\sigma}_{{\rm H}}$ $0.590(5)$ 0.191, 0.1840.1910.1840.191,\;0.1840.191 , 0.184
10) Gauss + BB Numeric PFC 0 $0.471(5)$ 0.172, 0.1560.1720.1560.172,\;0.1560.172 , 0.156
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.671}(5)$ 0.222, 0.1960.2220.1960.222,\;0.1960.222 , 0.196
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.675}(5)$ 0.212, 0.1920.2120.1920.212,\;0.1920.212 , 0.192
11) Gauss + BB Numeric Barlett 0 $0.466(5)$ 0.167, 0.1540.1670.1540.167,\;0.1540.167 , 0.154
$+5\bar{\sigma}_{{\rm H}}$ $0.633(5)$ 0.213, 0.1880.2130.1880.213,\;0.1880.213 , 0.188
$-5\bar{\sigma}_{{\rm H}}$ $0.635(5)$ 0.201, 0.1810.2010.1810.201,\;0.1810.201 , 0.181
12) Gauss + BB Numeric CH 0 ${\bf 0.675}(5)$ 0.235, 0.2200.2350.2200.235,\;0.2200.235 , 0.220
$+5\bar{\sigma}_{{\rm H}}$ $0.884(3)$ 0.345, 0.3520.3450.3520.345,\;0.3520.345 , 0.352
$-5\bar{\sigma}_{{\rm H}}$ $0.884(3)$ 0.347, 0.3540.3470.3540.347,\;0.3540.347 , 0.354
13) Gauss + BB Numeric FC Cheat 0 ${\bf 0.677}(5)$ 0.228, 0.2060.2280.2060.228,\;0.2060.228 , 0.206
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.682}(5)$ 0.232, 0.2100.2320.2100.232,\;0.2100.232 , 0.210
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.221, 0.1990.2210.1990.221,\;0.1990.221 , 0.199
14) Gauss + BB Numeric Heuristic 0 ${\bf 0.675}(5)$ 0.229, 0.2220.2290.2220.229,\;0.2220.229 , 0.222
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.703}(5)$ 0.241, 0.2320.2410.2320.241,\;0.2320.241 , 0.232
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.700}(5)$ 0.242, 0.2330.2420.2330.242,\;0.2330.242 , 0.233
Table 3: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
The outcome of the numerical study can be summarized as follows.
  1. 1.
    The Poisson ($2^{\rm nd}$ row) and Gaussian ($3^{\rm rd}$ row) likelihood provide consistent results in terms of coverage and $1\sigma$ uncertainty when using the definition of confidence intervals based on the Hessian matrix. In the Gaussian approximation, numerical minimization ($3^{\rm th}$ row) or an analytical solution ($4^{\rm th}$ row) provide identical results. As expected, by neglecting the stochastic nature of the MC templates, these likelihoods bring to severe under-coverage: the uncertainties are about twice smaller than $\bar{\sigma}_{\rm H}$ ($1^{\rm st}$ row).
  2. 2.
    The Hessian uncertainty for the Gauss + MC stat. likelihood ($5^{\rm th}$ row) gives the correct coverage for the nominal data set, but under-covers for the alternate data sets. The $1\sigma$ uncertainty is still smaller than $\bar{\sigma}_{\rm H}$, further suggesting that these confidence intervals will generally under-cover. Results obtained by using either the Gauss + MC stat. ($5^{\rm th}$ row) or Gauss + BBlite likelihood ($6^{\rm th}$ row) are quantitatively similar.
  3. 3.
    The Gauss + BB likelihood with Hessian uncertainty ($7^{\rm th}$ row) provides uncertainties that are finally larger than in the asymptomatic limit, but intervals still under-cover for both the nominal and alternate models. The level of under-coverage is similar for the three cases, suggesting that this method can at least provide confidence intervals with uniform coverage.
    It has been verified for several pseudo-data sets that a numerical minimization of the Poisson or Gaussian + BB likelihoods leads to identical numerical results. Motivated by the practical equivalence between the two, in the following we will concentrate on the latter since it can be minimized more efficiently.
  4. 4.
    Hessian ($7^{\rm th}$ row) or profile-likelihood ratio ($8^{\rm th}$ row) uncertainties are similar, showing that for each pseudo-data the log-likelihood function is reasonably parabolic around the minimum. However, the sampling distribution of $t_{\mu}$ is remarkably different from a chi-square with one degree of freedom, as illustrated by figure 2 for $\mu\equiv\mu_{1}^{\prime}=0$ on pseudo-data generated at the true value $\mu_{\rm t}=0$. The sampling distributions of the MLE estimator $\hat{\mu}$ and of the estimated Hessian uncertainty $\hat{\sigma}_{\rm H}$ are shown in figure 3. They both deviate visibly from a Gaussian distribution, with the latter featuring a long tail towards the right. Similar results are obtained when considering the other parameterization.
  5. 5.
    Fitting $(\mu_{1},\mu_{2})$ or $(\mu_{1}^{\prime},\mu_{2}^{\prime})$ leads to qualitatively very similar results in terms of coverage. Numerical differences for $\mu_{\rm t}=1\pm 5\bar{\sigma}_{\rm H}$ are mostly due to the different definition of alternate data, see e.g. eq. (4) and eq. (11). Hence, we conclude that under-coverage occurs independently from the parametrization being used.
Figure 2: Sampling distribution of the PLR function $t_{\mu}$ evaluated at $\mu\equiv\mu^{\prime}_{1}=0$ for data distributed according to the nominal model (i.e. with $\mu_{\rm t}=0$) with $N=2\times 10^{6}$, $\epsilon=0.03$, $n=200$, and $k=1$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments.
Figure 3: Sampling distributions of the point-estimator $\hat{\mu}\equiv\hat{\mu}^{\prime}_{1}$ and of its Hessian uncertainty $\hat{\sigma}_{\rm H}$ obtained from the Gauss + Barlow-Beeston likelihood applied to data distributed according to the nominal model (i.e. with $\mu_{\rm t}=0$) with $N=2\times 10^{6}$, $\epsilon=0.03$, $n=200$, and $k=1$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. For both distributions, a Gaussian fit is overlaid for reference.
The main conclusion of this study is that CI-setting methods based on the asymptotic properties of the MLE or PLR are clearly invalidated even though the sample sizes are large in an absolute sense for both data and simulation. This motivates us to go beyond the asymptotic approximation.
2.4 Beyond the asymptotic approximation
We have considered the following CI-setting methods inspired by the classical Feldman and Cousins (FC) unified approach FC . The main feature common to all methods is their use of the profile-likelihood ratio as an ordering principle with no priori assumption on its statistical distribution. The latter is indeed determined a posteriori from a collection of randomly generated pseudo-experiments. At variance with the original Feldman-Cousins formulation, the main challenge here is represented by the presence of nuisance parameters.
  • Profiled Feldman-Cousins
    If the asymptotic approximation is not valid, the profile-likelihood ratio $t_{\mu}$ in eq. (16) is not necessarily chi-square distributed and the coverage properties of eq. (17) are not guaranteed. In the so-called profiled Feldman-Cousins approach acero2024 , a best guess of the nuisance parameters $\hat{\boldsymbol{\theta}}_{\mu}$ is first obtained by minimizing the NLL function at a fixed value of $\mu$. Then, for each point $\mu$ sampling the range of values that we wish to include/exclude from the confidence interval, an ensemble of so-called Feldman-Cousins pseudo-experiments FC are generated by drawing event counts $y_{i}$ around $f_{i}(\mu,\hat{\boldsymbol{\theta}}_{\mu})$. The sampling distribution $P_{\rm FC}$ of $t_{\mu}$ is then used to determine a critical value $c_{\mu}(\alpha)$ at significance level $\alpha$, such that
[…]
    The confidence interval $I_{\rm FC}$ is defined as the set of points for which $t_{\mu}$ is less than the critical value $c_{\mu}(\alpha)$. Here, we are interested in understanding whether an interval based on eq. (19) would contain $\mu_{\rm t}$ with the right probability for at least the nominal and alternate toy models. For this purpose, we determine the critical values $c_{\mu}(\alpha)$ for $\mu_{k}\in\{1,1\pm 5\bar{\sigma}_{{\rm H}}\}$ and evaluate the coverage for pseudo-data distributed according to $\mu_{\rm t}=\mu_{k}$. Since we mostly care about coverage, we don’t perform the full Feldman-Cousins procedure to determine an exact confidence interval for each pseudo-experiment, which would be very time demanding, but follow instead a simplified procedure, namely, for the true and alternate models we define the confidence interval as
    $\{\mu\,:\,t_{\mu}\leq c_{\mu_{k}}\},\;\;\;{\rm with}\;\;\;\mu_{k}\in\{1,1\pm 5\bar{\sigma}_{{\rm H}}\}.$ (20)
    This construction leads to what we denote by profiled FC intervals. Notice that eq. (20) is not equivalent to the full Feldman-Cousins construction because the critical values $c_{\mu}$ are not computed for all possible values of $\mu$. The two definitions would eventually lead to the same intervals if the sampling distribution $P_{\rm FC}$ were independent from $\mu$.
  • Simplified Feldman-Cousins
    We considered an alternative construction of the FC interval where the best-fit estimators $(\hat{\mu},\hat{\boldsymbol{\theta}})$ are first determined for all parameters, including the POI; then, an ensemble of FC pseudo-experiments are generated by drawing Poisson distributed event counts around $f_{i}(\hat{\mu},\hat{\boldsymbol{\theta}})$; finally, the sampling distribution of $t_{\mu}$ and critical value $c_{\mu}$ are computed for $\mu=\hat{\mu}$. The resulting interval, which will be denoted as simplified FC, is then defined similarly to eq. (20) with $\mu_{k}=\hat{\mu}$.
  • Profile-likelihood-ratio with Barlett’s correction
    In case the asymptotic approximation is not valid, it has been suggested by Barlett Barlett ; CowanUnc to modify the profile-likelihood ratio $t_{\mu}$ such that its expectation value equals its asymptotic value of one (for one degree of freedom), that is $t_{\mu}\to t^{\prime}_{\mu}=t_{\mu}/\langle t_{\mu}\rangle$. Thus, starting from the sampling distributions obtained from the FC pseudo-experiments, $\langle t_{\mu}\rangle$ is determined for $\mu\in\{1,1\pm 5\bar{\sigma}_{{\rm H}}\}$, and the corresponding confidence interval defined as
    $\{\mu\,:\,t^{\prime}_{\mu}\leq 1\}.$ (21)
  • Cousins-Highlands
    In the profiled FC approach, all nuisance parameters, including those related to the MC statistical uncertainties, are fixed to their best-fit to the data, but their post-fit uncertainty is neglected when the FC pseudo-experiments are generated. This can be partly accounted for by following an hybrid Bayesian-frequentist approach first proposed by Cousins and Highland COUSINS1992331 . In the Cousins-Highland (CH) approach, the bin contents $y_{i}$ in each FC pseudo-experiment are drawn from different central values $f_{i}(\mu,\tilde{\boldsymbol{\theta}}_{\mu})$, where $\tilde{\boldsymbol{\theta}}_{\mu}$ is sampled from a multivariate normal distribution with mean $\hat{\boldsymbol{\theta}}_{\mu}$ and covariance matrix given by the a posteriori covariance matrix of $\hat{\boldsymbol{\theta}}_{\mu}$. The resulting sampling distribution $P_{\rm HC}$ differs from $P_{\rm FC}$, generally leading to more conservative intervals.
  • Feldman-Cousins “cheat”
    Finally, we considered an ideal, albeit unrealistic, construction of confidence intervals which is still based on the PLR, but with all parameters fixed to their true values for pseudo-experiments generation. In this case, the sampling distribution of $t_{\mu}$ from the FC pseudo-experiments matches the sampling distribution of the pseudo-data so that the coverage should come out right by construction. We will use this definition as a reference for the true $1\sigma$ uncertainty.
Using the numerical setup described in section 2.2, we studied the performances of the newly introduced CI-setting methods. All results are reported in tables 2 and 3. The main observations can be summarized as follows.
  1. 1.
    The profiled Feldman-Cousins approaches ($9^{\rm th}$ and $10^{\rm th}$ rows) provide intervals that under-cover in general, although less than from using the PLR. However, the coverage is found to depend on the value of $\mu_{\rm t}$. As expected, the profiled method ($10^{\rm th}$ row) performs better than the simplified one ($9^{\rm th}$ row), suggesting that the distribution $P_{\rm FC}(t_{\mu})$ depends on $\mu$.
  2. 2.
    The Barlett correction to the PLR method ($11^{\rm th}$ row) provides results similar to the profiled FC construction ($10^{\rm th}$ row), showing that the sampling distribution $P_{\rm FC}$ from the FC pseudo-experiments is reasonably close to a chi-square with one degree of freedom. This is also in agreement with the previous observation that intervals based on either the Hessian matrix or the PLR function are almost identical.
  3. 3.
    The Cousins-Highlands method ($12^{\rm th}$ row) leads to the largest uncertainty, but the intervals now over-cover, with a coverage level that depends on $\mu_{\rm t}$.
  4. 4.
    The Feldman-Cousins “cheat” method provides the right coverage by construction: the right CI size for $\mu$ would be about twice as large as the one obtained from the Hessian or PLR method on the full Barlow-Beeston likelihood, i.e. what one would have obtained based on the asymptotic properties of maximum-likelihood estimators.
The main conclusion of the study is that none of the realistic CI-setting methods considered here can provide the right coverage for the toy model: the asymptotic properties of the PLR are clearly invalidated by non-negligible corrections implied by Wilks’ theorem, which might have been unexpected given the large size of the data sample or the number of bins; on the other hand, methods that do not rely on asymptotic properties suffer from the problem of handling nuisance parameters, most critically, those related to statistical fluctuations of the MC templates.
However, one could reasonably expect that both issues will be solved in the limit of negligible MC statistical uncertainties, and the right coverage will eventually be restored. This can be achieved in multiple ways, the most obvious being by increasing the size of the simulated samples relative to the data or, if possible, by using larger samples for both. Reducing the number of bins might also help. Alternatively, we might conclude that the problem is related to some intrinsic feature of the model. We test these hypotheses by studying alternate model configurations in the next subsection.
2.5 Recovering the correct coverage
The results obtained by using alternate values for $N$, $\epsilon$, $n$, and $k$ are collected in appendix B. The main observations from this study can be summarized as follows.
  1. 1.
    Increasing the size of the MC samples. Tables 5 and 6 collect the results obtained by increasing the statistical power of the MC simulation relative to the data ($k$) by factors of $10$ and $40$, respectively. As one would have naively expected, the impact of the limited MC statistics becomes increasingly less relevant as the MC sample size grows compared to the data. For $k=10$, the correct behavior is still not fully achieved, while for $k=40$ all likelihood functions and CI setting methods provide roughly the same intervals and the right coverage is recovered for all cases. This shows that the asymptotic properties of the MLE can be ultimately recovered once the MC template fluctuations become sufficiently small.
  2. 2.
    Improving the signal-to-background separation. The results reported in tables 7 and 8 have been obtained by choosing $\epsilon=0.06$ and $\epsilon=0.12$, respectively. As the value of $\epsilon$ increases, the model becomes better behaved, as also reflected by a reduction in the condition number of the Hessian matrix. The coverage of the Barlow-Beeston likelihood with Hessian or PLR uncertainties improves, but the correct coverage cannot be recovered even for $\epsilon\sim{\cal O}(10\%)$. The level of under-coverage from the FC methods remains stable.
  3. 3.
    Reducing the number of bins. The results reported in tables 9 and 10 have been obtained by rebinning the histograms by factors of $2$ and $10$, respectively. As $n$ gets smaller, the coverage of the Hessian- or PLR-method improves steadily, but the intervals slightly under-cover even for $n=20$. As expected, fluctuations of MC templates across bins tend to average out with a coarser binning. Again, the level of under-coverage from the FC method does not change significantly.
  4. 4.
    Increasing the size of both data and simulation. The results in tables 11 and 12 have been obtained by increasing the number of events in data by factors of $10$ and $100$, respectively, while keeping the same proportion between data and simulation. For sufficiently large values of $N$, the asymptotic limit is ultimately reached when using the full Barlow-Beeston likelihood function (7), so that CI-setting methods based on the Hessian uncertainty or the PLR can finally provide the right coverage. The FC-methods still suffer from systematic under-coverage.
2.6 Discussion
We summarize the main results obtained from the toy study. The correct likelihood function for the problem under consideration is given by eq. (7): neglecting the stochastic nature of the MC templates $t_{ji}$, as done for example in eq. (6), has a sizable impact in terms of under-coverage already with an equal amount of data and simulated events. The “Barlow-Beeston lite” or “MC stat.” approaches treat the uncertain MC predictions as an additional source of statistical noise to the data. This accounts in part for the random fluctuations of MC templates, but is not sufficient to restore the right coverage, as demonstrated by the difference between using the lite vs. full Barlow-Beeston methods. When using the correct likelihood function, the asymptotic properties of maximum-likelihood estimators and of the profile-likelihood ratio are manifestly broken even if the number of events per bin is large enough to justify, for example, the approximation of the NLL as a chi-square function. We found that the asymptotic properties can eventually be recovered by applying a sufficiently large scaling of the data and/or simulation sizes, but we could not find a clear indication of which scale should be considered sufficiently “large”.
Approaches based on the post-fit distribution of the PLR suffer from a related problem: the post-fit nuisance parameters, in particular those related with the fluctuations of MC templates, are never exactly set to their true value. The footprint of the biased MC templates survives in the post-fit distribution of $t_{\mu}$, leading to an artificially augmented sensitivity to $\mu$. We remark that the problem of under-coverage of confidence intervals in the presence of nuisance parameters is well-known already conrad2002 ; acero2024 ; Punzi ; Cranmer . It can be ultimately ascribed to the fact that the standard Neyman’s construction for confidence intervals on one parameter of interest is not uniquely defined when the likelihood depends on multiple nuisance parameters, since some prescription on how to project a larger multi-dimensional space onto a one-dimensional closed interval must be supplied Punzi ; Cranmer . The profile-likelihood approach is just one of the infinitely many ways to achieve such a dimensional reduction. While this matter has been mostly addressed in low-statistics counting experiments and for relatively simple problems, we have shown that the problem can also arise in high-statistics measurements.
2.6.1 An heuristic confidence interval
Motivated by these considerations, we propose an heuristic confidence interval which combines Hessian uncertainties in the limit of infinite-size MC samples with those obtained from two variants of the Barlow-Beeston method. More precisely, we define
$I_{\rm heur}=\left[\hat{\mu}-\hat{\sigma}_{\rm heur},\;\hat{\mu}+\hat{\sigma}_{\rm heur}\right],$ (22)
where $\hat{\mu}$ is the maximum-likelihood estimator obtained by minimizing the full Barlow-Beeston likelihood from eq. (7), while $\hat{\sigma}_{\rm heur}$ is defined as
$\hat{\sigma}_{\rm heur}=\bar{\sigma}_{{\rm H}}\cdot\sqrt{1+\frac{1}{k}}\cdot\left(\frac{\hat{\sigma}_{\rm H}^{\rm BBfull}}{\hat{\sigma}_{\rm H}^{\rm BBlite}}\right),$ (23)
where $\bar{\sigma}_{{\rm H}}$ is the Hessian uncertainty in the limit of infinite MC statistics, $k$ is the statistical power of the MC simulation relative to the data, and $\hat{\sigma}_{\rm H}^{\rm BBfull}$ ($\hat{\sigma}_{\rm H}^{\rm BBlite}$) is the Hessian uncertainty derived from the full (lite) Barlow-Beeston likelihoods. We remark the importance of using the maximum-likelihood estimator $\hat{\mu}$ based on the full Barlow-Beeston likelihood rather than its lite approximation, since the latter does not provide an unbiased estimator of $\mu$.
The rationale behind eq. (23) is the following. The Hessian uncertainty derived from the full Barlow-Beeston likelihood underestimates the real statistical spread of $\hat{\mu}$ due to the presence of spurious constraints which break the asymptotic approximation. Conversely, the Hessian uncertainty $\bar{\sigma}_{{\rm H}}$ would match the standard deviation of $\hat{\mu}$, if the MC templates were exact. We take it as our staring point. Our intuition is that statistical uncertainties on the MC templates should at least add in quadrature to the Poisson uncertainty of the data, which corresponds to the square root in eq. (23). However, assigning an overall uncertainty to the total MC prediction per bin is not fully correct as it ignores the interplay between the $\mu$ parameter and MC statistical fluctuations pertaining to different templates. In fact, the existence of such an interplay is made evident by the breaking of the Barlow-Beeston lite approximation. The ratio between Barlow-Beeston full and lite Hessian uncertainties, corresponding to the last term on the right-hand side of eq. (23), provides an effective correction factor to account for the increased statistical uncertainty when the MC templates $\theta_{ji}$ are allowed to float in the fit independently from each other.
This heuristic definition ($14^{\rm th}$ row of tables 2 and 3) provides good coverage properties for all values of $\mu_{\rm t}$, as well as for different choices of model configurations. This suggests that $\hat{\sigma}_{\rm heur}$ could be at least used as a proxy for the real statistical spread of the $\hat{\mu}$ estimator in the presence of fluctuations of both the data and the simulated samples, and so it provides a practical test to gauge the goodness of asymptotic confidence intervals. This observation is intriguing and perhaps worth being further investigated, although a rigorous proof of its validity for a generic problem is beyond the scope of this work. We should remark, however, that $\hat{\sigma}_{\rm heur}$ is defined in terms of $\bar{\sigma}_{\rm H}$, which for a realistic problem might not be known a priori. In sec. 3.6 we will address the problem of estimating $\bar{\sigma}_{\rm H}$ when only finite-size MC samples are available.
Two remarks are worth making here. First, we note that eq. (23) might prove difficult to compute for realistic models that are more complex than our simple toy model, especially when the likelihood function has to handle correlations between different templates. In these cases, more sophisticated solutions should be found. Second, we have restricted our discussion to confidence intervals in one dimension, namely for a single parameter of interest, although in some cases one might be interested in setting confidence intervals in more dimensions. How to assess (and possibly correct) under-coverage in more dimensions is beyond the scope of this work.
In the remainder of the paper, we will study the problem of under-coverage of asymptotic confidence intervals from a more general perspective with the goal of identifying the origin of the observed effect.
3 The general case
The main scope of this section is to characterize in more generality the role played by statistical fluctuations of MC templates in the determination of confidence intervals constructed in the asymptotic approximation. For this purpose, the following approach is followed. We start by assuming that the MC templates are exact. This approximation allows us to derive a set of well-known analytic results for point estimators and confidence intervals based on the asymptotic properties of maximum-likelihood estimators Behnke . The same formalism can also accommodate extensions of the likelihood function that account for finite-size effects in an approximate way, such as the Gauss + MC stat. or Barlow-Beeston lite approaches as discussed in section 2. Then, we will treat statistical fluctuations of MC templates as perturbations on top of the true unknown templates and collect their contributions to asymptotic formulas as an overall bias term, whose sign and magnitude will then be the main subject of study.
3.1 Methodology
Consider a binned data set represented by independent Poisson-distributed event counts ${\bf y}$ with expectation value parametrized by the model function ${\bf f}(\mu,\boldsymbol{\theta})$, where $\mu$ is the parameter of interest and $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{p})$ is a vector of nuisance parameters, which are assumed to be not constrained by auxiliary measurements. The generalization to the case in which both constrained and unconstrained parameters are present is discussed in section 3.7.
In the limit of large event counts, that is $y_{i}\gg 1$ for $i=1,\ldots,n$, the negative log-likelihood function can be approximated by
$-\ln{L}(\mu,{\boldsymbol{\theta}})=\frac{1}{2}\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta})\right)^{T}{\bf V}^{-1}\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta})\right)+\ldots$ (24)
where ${\bf V}={\rm diag}({\bf y})$, and the dots stand for an overall constant plus terms that are of higher-order in $(y_{i}-f_{i})$. Equation (24) is the well-known Neyman’s $\chi^{2}$ test-statistic Neyman . We take ${\bf f}$ to be a differentiable function of $(\mu,\boldsymbol{\theta})$ and expand it around a suitable initial point $(\mu_{0},\boldsymbol{\theta}_{0})$:
${\bf f}(\mu,\boldsymbol{\theta})={\bf f}_{0}+{\bf j}(\mu-\mu_{0})+{\bf J}(\boldsymbol{\theta}-\boldsymbol{\theta}_{0})+\cdots,$ (25)
where ${\bf j}$ is the gradient of ${\bf f}$ with respect to $\mu$, ${\bf J}$ is the $n\times p$ Jacobian matrix of ${\bf f}$ with respect to ${\boldsymbol{\theta}}$, and the dots stand for higher-order terms in the Taylor expansion. While this approximation is exact for linear functions, such as the toy model of section 2, for generic functions it should be iterated around the intermediate minima of eq. (24), until convergence [^4] Clearly, the validity of the linear approximation is model dependent and should be assessed case by case. In particular, model functions that change in a very asymmetric way around the true value of the parameters (e.g. due to non-negligible contributions from the second-order derivatives) should be handled with care.. Hence, we will assume that ${\bf j}$ and ${\bf J}$ are the gradients and Jacobian matrix evaluated at a point in the parameter space which is sufficiently close to the global minimum of the NLL function. Without loss of generality, we take ${\bf J}$ as a full-rank matrix, otherwise the model should be redefined to remove the degeneracy.
3.1.1 The profile-likelihood ratio
Thanks to eq. (25), the function on the right-hand side of eq. (24) can be minimized by solving a linear system of equations, leading to the well-known linear least-square estimators Cowan . To this aim, we can first put the right-hand side of eq. (24) in the more evocative form
$-2\ln L(\mu,{\boldsymbol{\theta}})=\lVert{\bf r}_{\mu}-{\bf A}(\boldsymbol{\theta}-\boldsymbol{\theta}_{0})\rVert^{2},$ (26)
where the newly introduced quantities are defined as
${\bf r}_{\mu}={\bf d}-{\bf b}(\mu-\mu_{0})\;\;\;{\rm and}\;\;\;\left({\bf A},{\bf b},{\bf d}\right)\equiv{\bf V}^{-\frac{1}{2}}\left({\bf J},{\bf j},{\bf y}-{\bf f}_{0}\right),$ (27)
The point $\hat{\boldsymbol{\theta}}_{\mu}$ that minimizes eq. (26) at a fixed value of $\mu$ is given by
$\hat{\boldsymbol{\theta}}_{\mu}=\boldsymbol{\theta}_{0}+({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}{\bf r}_{\mu}.$ (28)
The matrix ${\bf W}\equiv{\bf A}^{T}{\bf A}$ is positive-definite, hence invertible, because ${\bf A}$ was assumed to have full rank. By inserting the right-hand side of eq. (28) into eq. (26), the log-likelihood function becomes a quadratic function of $\mu$, which can be again minimized analytically yielding the estimator:
$\hat{\mu}=\mu_{0}+\left({\bf b}^{T}{\bf U}{\bf b}\right)^{-1}{\bf b}^{T}{\bf U}{\bf d},$ (29)
with the matrix ${\bf U}$ defined as
${\bf U}={\bf 1}-{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}.$ (30)
After some straightforward algebraic manipulation, we can finally write the profile log-likelihood function as
$t_{\mu}=-2\ln\frac{L(\mu,\hat{\boldsymbol{\theta}}_{\mu})}{L(\hat{\mu},\hat{\boldsymbol{\theta}})}={\bf b}^{T}{\bf U}{\bf b}\cdot(\mu-\hat{\mu})^{2},$ (31)
where $\hat{\boldsymbol{\theta}}$ is the vector of nuisance parameters that minimizes eq. (24). According to Wilks’ theorem Wilks $t_{\mu}$ is asymptotically distributed as a chi-square with one degree-of-freedom, so that the quadratic form
$S={\bf b}^{T}{\bf U}{\bf b}={\bf b}^{T}{\bf b}-{\bf b}^{T}{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}{\bf b}$ (32)
converges to the reciprocal of the variance of $\hat{\mu}$. A consequence of eq. (32) is that $\mu$ cannot be measured (i.e. $\sigma_{\rm H}$ would be virtually infinite) when ${\bf b}\in{\rm Span}\{\bf A\}$ [^5] Physically, this would imply that a point in the NP space can be always found such that the right-hand side of eq. (26) is invariant under a variation of $\mu$.: by definition, in this case one can write ${\bf b}$ as a linear combination of the columns of ${\bf A}$, that is ${\bf b}={\bf A}{\bf w}$ for some vector of weights ${\bf w}$, so that
${\bf U}{\bf b}$ $=({\bf 1}-{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}){\bf A}{\bf w}={\bf A}{\bf w}-{\bf A}({\bf A}^{T}{\bf A})^{-1}({\bf A}^{T}{\bf A}){\bf w}$
$={\bf A}{\bf w}-{\bf A}{\bf w}={\bf 0},$ (33)
which then implies $S=0$.
An alternative way to prove the relation between the $S$ variable defined in eq. (32) and the variance of the MLE estimator $\hat{\mu}$ is to derive directly the full set of MLE estimators, that is the point in the parameter space that minimizes eq. (24) globally, obtaining:
$\begin{pmatrix}\hat{\boldsymbol{\theta}}\\ \hat{\mu}\end{pmatrix}=\begin{pmatrix}{\boldsymbol{\theta}}_{0}\\ {\mu}_{0}\end{pmatrix}+\left({\bf B}^{T}{\bf B}\right)^{-1}{\bf B}^{T}{\bf d},$ (34)
where ${\bf B}$ is a $n\times(p+1)$ matrix obtained by adding the column ${\bf j}$ on the right of ${\bf J}$, and ${\bf C}=\left({\bf B}^{T}{\bf B}\right)^{-1}$ is the covariance matrix of $(\hat{\boldsymbol{\theta}},\hat{\mu})$. By using the formula for the inverse block matrix, one can easily verify that
${\bf C}=\begin{pmatrix}{\bf W}&{\bf A}^{T}{\bf b}\\ {\bf b}^{T}{\bf A}&{\bf b}^{T}{\bf b}\end{pmatrix}^{-1}=\begin{pmatrix}{\bf W}^{-1}+S^{-1}{\bf W}^{-1}{\bf A}^{T}{\bf b}{\bf b}^{T}{\bf A}{\bf W}^{-1}&-S^{-1}{\bf W}^{-1}{\bf A}^{T}{\bf b}\\ -S^{-1}{\bf b}^{T}{\bf A}{\bf W}^{-1}&S^{-1}\end{pmatrix}.$ (35)
Equation (35) implies that
${\sigma}^{2}_{\rm H}=\left[{\bf C}\right]_{p+1,\,p+1}=S^{-1},$ (36)
thus showing that, in the Gaussian approximation, the standard deviation of $\hat{\mu}$ obtained from the Hessian matrix and from the concavity of the profile-likelihood ratio function coincide [^6] Formally, eq. (35) implies that $S$ is the Schur complement of the upper-left block ${\bf W}$ of the Hessian matrix ${\bf C}^{-1}$..
The matrix ${\bf U}$ introduced in eq. (30) is a $n\times n$ symmetric, positive semi-definite, and idempotent matrix, that is ${\bf U}^{T}{\bf U}={\bf U}$, and so is ${\bf 1}-{\bf U}$. Therefore, we can write it as
${\bf U}={\bf 1}-\sum_{j=1}^{p}{\bf u}_{j}{\bf u}_{j}^{T},$ (37)
where ${\bf u}_{j}$ are linearly independent and unit-norm vectors associated with the $p$ unit eigenvalues of ${\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}$. The ${\bf u}_{j}$ eigenvectors provide an orthonormal basis to ${\rm Span}\{{\bf A}\}$, i.e. the linear space spanned by its columns $\{{\bf a}_{1},\ldots,{\bf a}_{p}\}$. Being associated with identical eigenvalues, the vectors ${\bf u}_{j}$ obtained from the eigen-decomposition of ${\bf 1}-{\bf U}$ are not necessarily orthogonal, but can always be made such by using e.g. the Gram–Schmidt process.
Based on eq. (32), two limiting cases can be identified:
  • the vector ${\bf b}$ has a large projection onto ${\rm Span}\{\bf A\}$, that is
    ${\bf b}=\sum_{j=1}^{p}\left({\bf b}^{T}{\bf u}_{j}\right){\bf u}_{j}+{\bf b}_{\perp},$
    with ${\bf b}_{\perp}\notin{\rm Span}\{{\bf A}\}$ and $\lVert{\bf b}_{\perp}\rVert\ll\lVert{\bf b}\rVert$;
  • the vector ${\bf b}$ is almost orthogonal to ${\rm Span}\{{\bf A}\}$.
We see from eq. (32) that the extent to which a problem can be classified according to either of the two cases can be quantified by the positive-definite metric $\rho_{\mu}$ defined by
$\rho^{2}_{\mu}=\frac{{\bf b}^{T}{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}{\bf b}}{{\bf b}^{T}{\bf b}}\leq 1,$ (38)
[…]
$1-\left(\left[{\bf C}\right]_{p+1\,p+1}\cdot\left[{\bf C}^{-1}\right]_{p+1\,p+1}\right)^{-1},$ (39)
as defined for example in Minuit JAMES1975343 . Finally, we notice that $\rho^{2}_{\mu}{\sigma}^{2}_{\rm H}$ is also equal to the difference in quadrature between ${\sigma}^{2}_{\rm H}$ and the variance of $\mu$ when $\boldsymbol{\theta}$ are fixed to their post-fit values. Hence, $\rho_{\mu}$ can be in principle computed also from the group impact ($I_{\boldsymbol{\theta}}$) on $\mu$ from all nuisance parameters Pinto:2023yob ; CMS:2024lrd , namely:
$\rho_{\mu}=\frac{I_{\boldsymbol{\theta}}}{{\sigma}_{\rm H}},\;\;\;{\rm with}\;\;\;I_{\boldsymbol{\theta}}=\sqrt{\left[{\bf C}\right]_{\boldsymbol{\theta}\mu}^{T}\left[{\bf C}\right]_{\boldsymbol{\theta}\boldsymbol{\theta}}^{-1}\left[{\bf C}\right]_{\boldsymbol{\theta}\mu}},$ (40)
where $\left[{\bf C}\right]_{\boldsymbol{\theta}\boldsymbol{\theta}}$ and $\left[{\bf C}\right]_{\boldsymbol{\theta}\mu}$ stand for the upper-left $p\times p$ and upper-right $p\times 1$ blocks of the covariance matrix ${\bf C}$ [^7] For a single parameter $\theta_{j}$, the impact can be more simply written as $I_{\theta_{j}}=\frac{\left[{\bf C}\right]_{\mu\theta_{j}}}{\sqrt{\left[{\bf C}\right]_{\theta_{j}\theta_{j}}}}=|\rho_{\mu\theta_{j}}|\cdot\sigma_{\rm H}$, with $\rho_{\mu\theta_{j}}$ equal to the linear correlation between $\mu$ and $\theta_{j}$..
3.2 Impact of MC fluctuations on Hessian uncertainties
The results presented in section 3.1.1 assume that the model function ${\bf f}$ is known exactly. Suppose instead that ${\bf f}$ is determined from MC samples consisting of finite numbers of events, thus resulting in perturbed estimators $\hat{\bf b}={\bf b}+\boldsymbol{\beta}$ and $\hat{\bf A}$, with the latter giving rise to a perturbed matrix $\hat{\bf U}$, defined as in eq. (30) with the replacement ${\bf A}\to\hat{\bf A}$. Since $\hat{\bf U}$ must also be idempotent, it admits an expansion in terms of a new basis of eigenvectors $\hat{\bf u}_{j}={\bf u}_{j}+\boldsymbol{\nu}_{j}$. By expanding $S$ around its true value, we thus obtain an estimator $\hat{S}$ given by
$\hat{S}$ $=\hat{\bf b}^{T}\left(1-\sum_{j=1}^{p}\hat{\bf u}_{j}\hat{\bf u}_{j}^{T}\right)\hat{\bf b}={S}+2\boldsymbol{\beta}^{T}{\bf U}{\bf b}+\boldsymbol{\beta}^{T}{\bf U}\boldsymbol{\beta}$
$-2\sum_{j=1}^{p}\left({\bf b}^{T}{\bf u}_{j}\right)\left({\bf b}^{T}\boldsymbol{\nu}_{j}\right)-{\bf b}^{T}\left(\sum_{j=1}^{p}\boldsymbol{\nu}_{j}\boldsymbol{\nu}_{j}^{T}\right){\bf b}$
$+2\boldsymbol{\beta}^{T}\left(-\sum_{j=1}^{p}{\bf u}_{j}\boldsymbol{\nu}_{j}^{T}-\sum_{j=1}^{p}\boldsymbol{\nu}_{j}{\bf u}_{j}^{T}\right){\bf b}+\ldots,$ (41)
where the dots stand for terms that are of higher-order in the perturbations. We now take the expectation value of the right-hand side of eq. (3.2) in the space of MC predictions, that is the mean value of $\hat{S}$ if infinitely many identical replicas of the MC templates could be tested, obtaining:
$\langle\hat{S}\rangle$ $={S}+2\langle\boldsymbol{\beta}^{T}\rangle{\bf U}{\bf b}+{\rm Tr}\left[{\bf U}\langle\boldsymbol{\beta}\boldsymbol{\beta}^{T}\rangle\right]$
$-2\sum_{j=1}^{p}\left({\bf b}^{T}{\bf u}_{j}\right)\left({\bf b}^{T}\langle\boldsymbol{\nu}_{j}\rangle\right)-{\bf b}^{T}\left(\sum_{j=1}^{p}\langle\boldsymbol{\nu}_{j}\boldsymbol{\nu}_{j}^{T}\rangle\right){\bf b}$
$-2\sum_{j=1}^{p}\left({\rm Tr}\left[\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle\,{\bf u}_{j}{\bf b}^{T}\right]+{\rm Tr}\left[\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle\right]{\rm Tr}\left[{\bf u}_{j}{\bf b}^{T}\right]\right)+\ldots,$ (42)
The correlation terms on the last line of eq. (3.2) are model dependent, so additional assumptions are needed to proceed. We notice that in many cases the gradient vectors of ${\bf f}$ with respect to its $p+1$ parameters can be computed independently from each other. For example, this applies when $\mu$ and $\boldsymbol{\theta}$ are strength modifiers relative to templates obtained from independently simulated processes, which is, for example, the case considered by Barlow and Beeston in their original work BARLOW1993219 . Conversely, there is at least another relevant case in which the assumption of uncorrelated templates might seem inappropriate, that is when nuisance parameters are associated with systematic uncertainties estimated from event-by-event re-weightings of the “nominal” MC sample. In this case, correlations between the columns of ${\bf A}$ cannot be ignored a priori because of correlated Poisson fluctuations of the MC sample. However, if the dominant source of statistical uncertainty on the MC prediction originates from the variance of the event weights rather than from the purely Poisson uncertainty, different re-weightings of the same MC sample could be still treated as approximately independent. To proceed further, we make the assumption that the covariance matrix $\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle$ is zero and consider the remaining terms.
3.3 Contributions from the Jacobian matrix
Let us start by analyzing the ${\bf U}$-dependent bias terms appearing in eq. (3.2). The expectation values $\langle\boldsymbol{\nu}_{j}\rangle$ are generally different from zero. Indeed, from the condition $\lVert\hat{\bf u}_{j}\rVert=1$ it follows that
$02{\bf u}_{j}^{T}\langle\boldsymbol{\nu}_{j}\rangle=-\langle\boldsymbol{\nu}_{j}^{2}\rangle<$ 0 . (43)
We notice that the effect of statistical fluctuations on the elements of $\hat{\bf A}$ is to stretch and rotate its column vectors. However, $S$ cannot depend on their normalization, since a global re-scaling of the columns can always be absorbed by a redefinition of their associated nuisance parameter [^8] This can be also proved by noticing that the matrix ${\bf U}$ is invariant under a right multiplication of ${\bf A}$ by an invertible matrix $\boldsymbol{\Lambda}$, that is ${\bf A}\to{\bf A}\boldsymbol{\Lambda}$, of which a column-wise re-scaling is a special case.. Without loss of generality, we can then assume $\lVert{\bf a}_{j}\rVert=1$ and fix their orientation by requiring e.g. ${\bf a}_{j}^{T}{\bf b}\geq 0$. Random fluctuations on the normalized vectors ${\bf a}_{j}$ generate perturbations $\boldsymbol{\alpha}_{j}$ corresponding to small rotations in a $n$-dimensional space, which must however average to zero when projected onto the sub-space orthogonal to ${\bf a}_{j}$, so that:
$\langle\boldsymbol{\alpha}_{j}\rangle=-\frac{\langle\boldsymbol{\alpha}_{j}^{2}\rangle}{2}{\bf a}_{j},$ (44)
where ${\langle\boldsymbol{\alpha}_{j}^{2}\rangle}$ is the variance of the angle by which the vector ${\bf a}_{j}$ is rotated. It is easy to see that ${\langle\boldsymbol{\alpha}_{j}^{2}\rangle}$ is of order $n\cdot\Delta_{j}^{2}$, with $\Delta_{j}$ being the average relative MC uncertainty on the elements of ${\bf a}_{j}$. Under the same assumptions, one can easily prove that $\langle\boldsymbol{\alpha}_{j}\boldsymbol{\alpha}_{j}^{T}\rangle$ must be a matrix that projects onto the subspace orthogonal to ${\bf a}_{j}$ with equal probability along each of its $n-1$ directions, that is
$\langle\boldsymbol{\alpha}_{j}\boldsymbol{\alpha}_{j}^{T}\rangle=\frac{\langle\boldsymbol{\alpha}^{2}_{j}\rangle}{n-1}\left({\bf 1}-{\bf a}_{j}{\bf a}_{j}^{T}\right).$ (45)
By means of eq. (44) and (45), the right-hand side of eq. (3.2) can be expanded further.
3.3.1 Orthogonal column vectors
We start by considering the case ${\bf a}_{j}^{T}{\bf a}_{k}=\delta_{jk}$, that is the column vectors of the Jacobian matrix ${\bf A}$ are orthogonal to each other, which also implies a diagonal matrix ${\bf W}$. This hypothesis makes the analytic treatment simpler and its interpretation more intuitive, without precluding the general validity, as discussed later. It also has some physical interest: for example, it is appropriate when nuisance parameters parametrize systematic effects that behave in an almost “orthogonal” way in the space of observables ${\bf y}$, for example when they are determined from a PCA analysis Pearson01111901 ; Hotelling . Under the assumption of orthonormal column vectors, we can exploit the arbitrariness in the definition of the ${\bf u}_{j}$ basis by choosing ${\bf u}_{j}={\bf a}_{j}$. The second line on eq. (3.2) can be expanded by means of eq. (44) and (45), obtaining after some straightforward calculation:
$\frac{n}{n-1}\sum_{j=1}^{p}\left({\bf b}^{T}{\bf a}_{j}\right)^{2}\langle\boldsymbol{\alpha}_{i}^{2}\rangle-\frac{{\bf b}^{T}{\bf b}}{n-1}\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right)$ (46)
Let us first consider the case of large global correlation, that is we let $\rho_{\mu}\to 1$. Neglecting for a moment the numerical pre-factors, the two pieces appearing in eq. (46) are of the same order of magnitude, but of opposite sign. However, the former is scaled by a factor of $n$, thus giving a dominant positive contribution when the number of bins is large, which is our case of interest. This can be intuitively understood by noticing that when $n\gg p$, statistical fluctuations on ${\bf A}$ have a larger probability of shuffling around the basis vectors by introducing new directions along which ${\bf b}$ can have a finite projection, rather than rotate the vectors inside the same space, which would leave $S$ unchanged. These extra contributions add quadratically to $\lVert{\bf b}_{\perp}\rVert^{2}$ and can be numerically relevant if the latter is truly small. The same conclusion can be proved for generic $n$ under additional assumptions. Indeed, by expanding ${\bf b}$ as a linear combination of the vectors ${\bf a}_{j}$, and neglecting the ${\bf b}_{\perp}$ term, we have
$\frac{n}{n-1}\sum_{j=1}^{p}\left({\bf b}^{T}{\bf a}_{j}\right)^{2}\langle\boldsymbol{\alpha}_{i}^{2}\rangle-\frac{{\bf b}^{T}{\bf b}}{n-1}\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right)>$
$\frac{1}{n-1}\left[p\sum_{j=1}^{p}\left({\bf b}^{T}{\bf a}_{j}\right)^{2}\langle\boldsymbol{\alpha}_{i}^{2}\rangle-\left(\sum_{j=1}^{p}\left({\bf b}^{T}{\bf a}_{j}\right)^{2}\right)\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right)\right]$ (47)
where we have used the fact that $n>p$. The quantity within square brackets is positive-definite for all $n$ as for Chebyshev’s sum inequality, provided that ${\bf b}^{T}{\bf a}_{1}\geq\ldots\geq{\bf b}^{T}{\bf a}_{p}$ and $\langle\boldsymbol{\alpha}_{1}^{2}\rangle\geq\ldots\geq\langle\boldsymbol{\alpha}_{p}^{2}\rangle$, which includes both the case of uniform uncertainties on the columns, as well as the case of increasingly larger uncertainty for column vectors that have larger overlap with ${\bf b}$, which is also an intuitive result.
We now consider the other limiting case, that is $\rho_{\mu}\to 0$. The first term on the second line of eq. (46) can be now neglected compared to the second term, which is instead always negative. Thus, we see that in this case the effect of random fluctuations of the MC templates is to reduce the value of $\langle\hat{S}\rangle$ relative to ${S}$ by a factor of
$1-\frac{1}{n-1}\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right)$ (48)
This result is also in line with our intuition: if ${\bf b}$ is already orthogonal to the linear space spanned by the columns of ${\bf A}$, fluctuations on the latter can only generate spurious components of ${\bf b}$ in ${\rm Span}\{{\bf A}\}$, thus reducing the norm of ${\bf b}_{\perp}$. However, we notice that the relative decrease implied by eq. (48) is only of order of the squared relative uncertainty on the elements of $\tilde{\bf A}$, and is not scaled by $n$, so its overall impact might well be numerically less relevant than the effect discussed in the previous case.
Equations (46) and (48) can be verified, for example, in a simple two-dimensional setup, namely for $n=2$ and $p=1$. Modulo an overall normalization factor to the gradient vector and to the single-column Jacobian matrix, we can define
${\bf b}=\left(\begin{matrix}1\\ 0\end{matrix}\right)\;\;\;{\rm and}\;\;\;{\bf A}=\begin{pmatrix}s_{\gamma}\\ c_{\gamma}\end{pmatrix}\Rightarrow{\bf U}=\begin{pmatrix}c^{2}_{\gamma}&1-s_{\gamma}c_{\gamma}\\ 1-s_{\gamma}c_{\gamma}&s^{2}_{\gamma}\end{pmatrix},$ (49)
where $s_{\gamma}\equiv\sin\gamma$, $c_{\gamma}\equiv\cos\gamma$, with $\gamma\in[0,\pi]$ an angle that parametrizes the degree of correlation between $\mu$ and the nuisance parameter $\theta$, as illustrated by the left panel of figure 4. From eq. (49) we have that $S=U_{11}=c^{2}_{\gamma}$ and $\rho^{2}_{\mu}=s^{2}_{\gamma}=1-S$. We treat statistical fluctuations on ${\bf A}$ as a pair of independent and Gaussian-distributed random variables $(\alpha_{1},\alpha_{2})$, which, for simplicity, are assumed to have the same variance $\langle\alpha^{2}\rangle$. Under fluctuations, the matrix ${\bf U}$ is then changed to
${\bf U}\to\hat{\bf U}=\begin{pmatrix}\frac{(c_{\gamma}+\alpha_{2})^{2}}{(s_{\gamma}+\alpha_{1})^{2}+(c_{\gamma}+\alpha_{2})^{2}}&\frac{(s_{\gamma}+\alpha_{1}-c_{\gamma}-\alpha_{2})^{2}+(s_{\gamma}+\alpha_{1})(c_{\gamma}+\alpha_{2})}{(s_{\gamma}+\alpha_{1})^{2}+(c_{\gamma}+\alpha_{2})^{2}}\\ -&\frac{(s_{\gamma}+\alpha_{1})^{2}}{(s_{\gamma}+\alpha_{1})^{2}+(c_{\gamma}+\alpha_{2})^{2}}\end{pmatrix}$ (50)
Using eq. (50) evaluated at randomly generated points $(\alpha_{1},\alpha_{2})$, the estimator $\hat{S}=\hat{U}_{11}$ can be calculated and its mean value $\langle\hat{S}\rangle$ extracted from a large number of random samplings. The results are shown on the right of figure 4 as a function of the global correlation $\rho^{2}_{\mu}$ and for three representative values of $\langle\alpha^{2}\rangle$. For small values of $\rho^{2}_{\mu}$, the ratio $\langle\hat{S}\rangle/S$ is slightly less than one, i.e. the bias on $\hat{S}$ is negative, in agreement with eq. (48). Conversely, when $\rho^{2}_{\mu}$ increases, the ratio becomes larger than one (hence, the bias gets positive), eventually growing as $\langle\alpha^{2}\rangle/(1-\rho^{2}_{\mu})$ for $\rho^{2}_{\mu}\to 1$, in agreement with eq. (46).
Figure 4: Left: sketch of the two-dimenensional problem of eq. (49). Right: the ratio between the expectation value $\langle\hat{S}\rangle$ and the true value $S$ for the two-dimensional problem of eq. (49) as a function of the squared correlation coefficient $\rho^{2}_{\mu}$, and for three representative values of $\langle\alpha^{2}\rangle$.
3.3.2 Generic column vectors
Finally, we consider the more general case for the columns vectors ${\bf a}_{j}$. In this case, we can always find a set of $p$ orthonormal vectors such that
${\bf u}_{j}=\sum_{k=1}^{p}w_{jk}{\bf a}_{k}$ (51)
where $w_{jk}$ are the elements of a non-singular square matrix. After some straightforward algebra, the expectation value of the two last terms on the second line of eq. (3.2) can be written as:
$\sum_{j,k,l=1}^{p}w_{jk}w_{jl}\left({\bf b}^{T}{\bf a}_{k}\right)\left({\bf b}^{T}{\bf a}_{l}\right)\langle\boldsymbol{\alpha}_{l}^{2}\rangle\left(1+\frac{\delta_{kl}}{n-1}\right)-\frac{{\bf b}^{T}{\bf b}}{n-1}\sum_{j,k=1}^{p}w^{2}_{jk}\langle\boldsymbol{\alpha}_{k}^{2}\rangle$ (52)
Again, we first focus on the case of large overlap between ${\bf b}$ and ${\rm Span}\{\bf A\}$, which allows us to consider only the first term on the right-hand side of eq. (52) since the remaining terms are relatively suppressed by a factor of $1/n$. If we denote by ${\bf v}$ a vector with $k^{\rm th}$-element equal to ${\bf b}^{T}{\bf a}_{k}$, we see that the triple sum of eq. (52) can be written as
${\bf v}^{T}\left(\sum_{j=1}^{p}{\bf w}_{j}{\bf w}_{j}^{T}{\rm diag}(\langle\boldsymbol{\alpha}_{1}^{2}\rangle,\ldots,\langle\boldsymbol{\alpha}_{p}^{2}\rangle)\right){\bf v}$ (53)
The matrix defined between parentheses in eq. (53) is positive (semi-)definite if $\langle\boldsymbol{\alpha}_{k}^{2}\rangle>0$ for all (at least one) index $k$, so that the quadratic form cannot be negative. In the limiting case of small $\rho_{\mu}$, the rightmost term in eq. (52) dominates, and the same conclusions derived for the case of orthogonal columns holds.
For intermediate values of the global correlation, the existence of competition between positive and negative bias terms to $\langle\hat{S}\rangle$ does not allow for a conclusive statement on the overall sign, and the bias term should be determined case by case.
3.4 Contribution from the gradient
Finally, we consider the role of statistical fluctuations on the gradient vector. Since the MC prediction for $\hat{\bf b}$ is randomly distributed around the true value ${\bf b}$, the expectation value of $\boldsymbol{\beta}$ is zero, so that terms linear in $\langle\boldsymbol{\beta}\rangle$ do not contribute to the bias term in eq. (3.2). The quadratic term on the first line of eq. (46) is positive-definite. This can be also seen as consequence of Jensen’s inequality applied to $\hat{S}$, which is a convex function of $\hat{\bf b}$. If we make the reasonable assumption $\langle\beta_{i}\beta_{j}\rangle=\langle\beta^{2}_{i}\rangle\delta_{ij}$, that is the MC predictions for different bins are statistically uncorrelated, the quadratic term can be further expanded as
${\rm Tr}\left[{\bf U}\,\langle\boldsymbol{\beta}\boldsymbol{\beta}^{T}\rangle\right]=n\left(\sum_{i=1}^{n}\frac{1}{n}\langle\beta_{i}^{2}\rangle\right)-\sum_{j=1}^{p}\left(\sum_{i=1}^{n}u_{j,i}^{2}\langle\beta_{i}^{2}\rangle\right),$ (54)
where the two terms in parentheses correspond to different ways of averaging the bin-by-bin variances $\langle\beta_{i}^{2}\rangle$. If the distribution of the latter is narrowly concentrated around an average value $\overline{\langle\beta^{2}\rangle}$, the right-hand side of eq. (54) is of order $\sim(n-p)\overline{\langle\beta^{2}\rangle}$, so it grows linearly with the number of bins. Thus, we see that statistical fluctuations on ${\bf b}$ can induce sizable positive biases to $\langle\hat{S}\rangle$ regardless of the global correlation coefficient, provided that $(n-p)\overline{\langle\beta^{2}\rangle}$ is comparable to $S$.
3.5 Discussion and connection with the toy model
We now summarize the assumptions and the main implications of eq. (3.2). We have assumed that the MC estimators for ${\bf b}$ and ${\bf A}$ are independent from each other and that statistical fluctuations of their elements can be treated as Gaussian-distributed noise around their true values. We have considered separately the impact of MC fluctuations on the Jacobian matrix ${\bf A}$ and on the gradient vector ${\bf b}$ in the form of a bias term to the quadratic form $S$, defined by eq. (32). If ${\bf b}$ has a large overlap with the linear space spanned by the columns of ${\bf A}$, or, equivalently, when the global correlation coefficient (38) is sizable, then the bias terms associated with fluctuations of the Jacobian matrix are positive-definite and increase with the number of bins $n$. Fluctuations of the gradient vector ${\bf b}$ also grow with $n$ and are always positive-definite. In both cases, $\hat{S}$ would differ from the true value $S$ by a positive offset, implying that the Hessian uncertainty, estimated as $\hat{\sigma}_{\rm H}=\hat{S}^{-1/2}$, would be underestimated [^9] More correctly, a positive bias on $\hat{S}$ would also imply a negative bias on $\hat{\sigma}_{\rm H}$ if the variance of $\hat{S}$ is small, so that $\langle\hat{S}^{-1/2}\rangle\approx\langle\hat{S}\rangle^{-1/2}$. In particular, by Taylor expansion, we find that $\frac{3}{8}\frac{\langle\hat{S}^{2}\rangle-\langle\hat{S}\rangle^{2}}{\langle\hat{S}\rangle^{2}}\ll 1$ should hold, a condition that is more likely to be satisfied when the number of bins is large, as for the law of large numbers..
These findings can finally explain the observations made in the context of the toy model of sec. 2: when the original parametrization is considered (i.e. $\mu=\mu_{1}$), the linear correlation $\rho=\rho_{\mu}$ is indeed close to unity, implying a large bias term from fluctuations of the Jacobian matrix; for $\mu=\mu_{1}^{\prime}$, instead, the smaller correlation reduces the impact of a noisy Jacobian matrix, and the bulk of finite-size MC effects are now propagated to $\hat{\sigma}_{\rm H}$ through a noisy gradient vector $T_{1i}^{\prime}$.
Remarkably, none of the bias terms appearing in eq. (46) is directly proportional to the true value $S$ by a naive factor of ${\cal O}(1)\cdot k^{-1}$, with $k$ related to the effective statistical power of the simulation relative to the data. The lack of proportionality also means that the bias and true value can stand in a ratio which is a priori arbitrary. This is one of the main observations of this work. Incidentally, this also explains how the Barlow-Beeston lite approach BBlite can be insufficient to account for MC uncertainties. As discussed in sec. 13, the lite version of the Barlow-Beeston likelihood is practically equivalent to a rescaling of the Poisson variance of the data by a factor of $(1+k^{-1})$, implying an increase of the Hessian uncertainty $\hat{\sigma}_{\rm H}$ by the square root of the same factor. However, a bias in $\hat{S}$, which is both positive and comparable to $S$, might reduce $\hat{\sigma}_{\rm H}$ by a potentially larger factor, ultimately leading to the non-physical situation of confidence intervals that are smaller than they would be in the limit of infinite-size MC samples.
As a last step in our discussion, we address the question of providing quantitative estimates for the bias terms appearing in eq. (3.2).
3.6 Asymptotic uncertainties from finite-size samples
The bias term generated by statistical fluctuations of the gradient vector and Jacobian matrix could be estimated in principle from eq. (3.2) by computing the covariance matrices of ${\bf b}$ and of the column vectors of ${{\bf A}}$, for example by bootstrap resampling, and using their nominal MC predictions $\hat{\bf b}$ and $\hat{\bf A}$ in place of their true, unknown values. The asymptotic Hessian uncertainty $\bar{\sigma}_{\rm H}={S}^{-1/2}$, i.e. the Hessian uncertainty in the limit of infinite MC statistics, could be determined from eq. (3.2) after subtracting the bias contribution.
A possibly simpler way to estimate $\bar{\sigma}_{\rm H}$ consists in analyzing the scaling of the Hessian uncertainty $\hat{\sigma}_{\rm H}$ as a function of the MC-to-data ratio $k$ at a fixed size $N$ of the data. Indeed, the following scaling laws can be easily derived:
${\bf b}\sim\sqrt{N},\;\;\;{\bf U}\sim 1,\;\;\;\langle\boldsymbol{\alpha}_{j}^{2}\rangle\sim\frac{1}{kN},\;\;\;\langle\boldsymbol{\beta}\boldsymbol{\beta}^{T}\rangle\sim\frac{1}{k},$ (55)
from which one gets:
$\bar{\sigma}_{\rm H}=\hat{\sigma}_{{\rm H}}\cdot\sqrt{1+\frac{\Delta_{N}}{k}},$ (56)
where $\Delta_{N}$ is a constant independent from $k$, and $\hat{\sigma}_{{\rm H}}$ is the Hessian uncertainty obtained from a MC sample with $k$-times the statistical power of the data. The right-hand side of eq. (56) contains two unknowns: $\bar{\sigma}_{\rm H}$ and $\Delta_{N}$. If $N$ is kept fixed, but a fraction $1/\xi$, with $\xi>1$ of the MC events is randomly selected and used to construct a new model prediction $\hat{\bf f}$, then the new Hessian uncertainty will be a fraction $r_{\xi}<1$ of the uncertainty from the full sample. Then, one can readily solve for $\bar{\sigma}_{\rm H}$, obtaining:
$\bar{\sigma}_{\rm H}=\left(\frac{r_{\xi}\sqrt{\xi-1}}{\sqrt{\xi r^{2}_{\xi}-1}}\right)\hat{\sigma}_{{\rm H}}.$ (57)
For example, for the toy model of section 2, the mean Hessian uncertainty for $N=2\times 10^{6}$, $n=200$, and $\alpha=0.03$ is found to be $\hat{\sigma}_{{\rm H}}=0.0572$ ($0.0447$) when $k=1$ ($k=0.5$), from which the asymptotic uncertainty can be predicted to be $0.0956$ from eq. (57), in good agreement with the true value of $0.0950$.
Equation (57) assumes that the likelihood does not contain any explicit term to account for the statistical uncertainties on the MC templates. It can be easily extended to the case where the MC statistical uncertainties are treated à la Barlow-Beeston lite by exploiting the Gauss + MC stat. approximation (see sec. 2.1.2), thus modifying the scaling of ${\bf b}$ via an additional multiplicative factor $\left(1+\frac{1}{k}\right)^{-{1}/{2}}$, yielding:
$\bar{\sigma}_{\rm H}=\left(\frac{r_{\xi}\sqrt{\xi-1}}{\sqrt{\alpha(1+\frac{1}{k})r^{2}_{\xi}-(1+\frac{\xi}{k})}}\right)\hat{\sigma}_{{\rm H}}.$ (58)
Equations (57) and (58) could be improved by taking the expectation values of $\hat{\sigma}_{{\rm H}}$ over several bootstrap resamplings of the MC sample.
Although $\bar{\sigma}_{\rm H}$ does not provide a reliable estimate of the actual statistical spread of $\hat{\mu}$ induced by fluctuations of both the data and the MC samples, it can provide useful information nonetheless. For example, values of $\bar{\sigma}_{\rm H}$ significantly larger than $\hat{\sigma}_{\rm H}$ would signal the presence of spurious constraints on the POI which are likely to invalidate the asymptotic properties of maximum-likelihood estimators even in the presence of a proper account of MC statistical uncertainties, as done for example in the full Barlow-Beeston method. Moreover, $\bar{\sigma}_{\rm H}$ is one of the inputs needed to construct the heuristic confidence interval of eq. (23).
3.6.1 Lite vs full Barlow-Beeston
Suppose that $\bar{\sigma}_{\rm H}$ calculated from eq. (56) is in good agreement with $\hat{\sigma}_{\rm H}$, or, equivalently, that the Hessian uncertainty is stable when the size of the MC sample is reduced. One might still want to ensure that the lite version of the Barlow-Beeston method represents an adequate approximation for a given problem. For this purpose, we can use the formalism developed in the previous section to elaborate the following analysis. We start by noticing that the MLE estimator $(\hat{\boldsymbol{\theta}},\hat{\mu})$ can be seen as a function of ${\bf B}$, ${\bf f}_{0}$, and ${\bf y}$, which are all random variables of known variance. By applying the chain rule for differentiation, the gradient of $(\hat{\boldsymbol{\theta}},\hat{\mu})$ with respect to $B_{ik}$ can be computed as
$\frac{\partial}{\partial B_{ik}}\begin{pmatrix}\hat{\boldsymbol{\theta}}\\ \hat{\mu}\end{pmatrix}=\left(\frac{\partial\left({\bf B}^{T}{\bf B}\right)^{-1}}{\partial B_{ik}}\right){\bf B}^{T}{\bf d}+\left({\bf B}^{T}{\bf B}\right)^{-1}\left(\frac{\partial{\bf B}^{T}}{\partial B_{ik}}\right){\bf d}$
$=-\left({\bf B}^{T}{\bf B}\right)^{-1}\frac{\partial\left({\bf B}^{T}{\bf B}\right)}{\partial B_{ik}}\left({\bf B}^{T}{\bf B}\right)^{-1}{\bf B}^{T}{\bf d}+\left({\bf B}^{T}{\bf B}\right)^{-1}\left(\frac{\partial{\bf B}^{T}}{\partial B_{ik}}\right){\bf d}$
$=-\left({\bf B}^{T}{\bf B}\right)^{-1}\left(\boldsymbol{\Pi}^{T}_{ik}{\bf B}+{\bf B}^{T}\boldsymbol{\Pi}_{ik}\right)\begin{pmatrix}\hat{\boldsymbol{\theta}}\\ \hat{\mu}\end{pmatrix}+\left({\bf B}^{T}{\bf B}\right)^{-1}{\boldsymbol{\Pi}^{T}_{ik}}{\bf d}$
$={\bf C}\left(\boldsymbol{\Pi}^{T}_{ik}\left({\bf d}-{\bf B}\begin{pmatrix}\hat{\boldsymbol{\theta}}\\ \hat{\mu}\end{pmatrix}\right)-{\bf B}^{T}\boldsymbol{\Pi}_{ik}\begin{pmatrix}\hat{\boldsymbol{\theta}}\\ \hat{\mu}\end{pmatrix}\right),$ (59)
where $\boldsymbol{\Pi}_{ik}={\bf e}_{i}{\bf e}_{k}^{T}$, with ${\bf e}_{l}$ the $l^{\rm th}$ vector of the canonical basis. Calculating the gradient with respect to ${\bf f}_{0}$ is straightforward. If the MC predictions for $B_{ik}$ and $f_{0,i}$ can be treated as independent, Gaussian-distributed variables with standard deviations $\Delta B_{ik}$ and $\Delta f_{0,i}$, the contribution to the variance of the MLE estimator $\hat{\mu}$ induced by fluctuations of the MC templates can be determined from standard error propagation as
$\hat{\sigma}^{2}_{\rm MC}=\sum_{i=1}^{n}\left[\sum_{k=1}^{p}\left(\frac{\partial\hat{\mu}}{\partial B_{ik}}\right)^{2}(\Delta B_{ik})^{2}+\left(\sum_{k=1}^{p+1}{C}_{p+1\,k}B_{ik}\right)^{2}(\Delta f_{0,i})^{2}\right].$ (60)
Equation (60) can be easily extended to generic covariances for ${\bf B}$ and ${\bf f}_{0}$. However, it should be noticed that the right-hand side of eq. (60) can only be evaluated using the MC prediction $\hat{\bf B}$ and $\hat{\bf f}_{0}$ derived from finite-size MC samples, so eq. (60) might be a biased estimator of the real variance of $\hat{\mu}$ under fluctuations of the MC templates. With this caveat in mind, we could still estimate the total variance of the maximum-likelihood estimator $\hat{\mu}$ under fluctuations of both the data and the MC samples as
$\hat{\sigma}_{\rm H,\,tot}\approx\sqrt{\hat{\sigma}_{\rm H}^{2}+\hat{\sigma}_{\rm MC}^{2}}.$ (61)
Values of $\hat{\sigma}_{\rm MC}$ larger than the Barlow-Beeston lite expectation of $\hat{\sigma}_{\rm H}/\sqrt{k}$, or, equivalently values of $\hat{\sigma}_{\rm H,\,tot}$ in excess of the naive expectation $\hat{\sigma}_{\rm H}\sqrt{1+k^{-1}}$, should be taken as an indication that the lite approach is insufficient to provide a proper account of MC statistical uncertainties, suggesting that the full approach should be adopted instead. For example, for the toy study presented in section 2, we find that $\hat{\sigma}_{\rm MC}$ estimated from eq. (60) is a factor of $\sim 1.5$ larger than $\hat{\sigma}_{\rm H}/\sqrt{k}$ for the nominal model configuration. Interestingly, this ratio decreases monotonically towards $\sim 1$ when $N$ or $\epsilon$ are increased, or when $n$ is reduced, corroborating the observation that the Barlow-Beeston lite approach becomes more efficient in these limits. Conversely, we find that the ratio approaches a plateau of $\sim 1.6$ for $k\gg 1$, when $n$, $N$ and $\epsilon$ are fixed to their default values, implying that the Barlow-Beeston lite version will never be equivalent to its full implementation for at least the nominal model configuration [^10] In practice this difference is irrelevant when $k$ is large because the contribution from MC statistical uncertainties to the total variance will also be negligible in this limit.. A summary of numerical results is reported in table 4.
$k\times(1,10,40)$ $n\times(\frac{1}{2},\frac{1}{10},\frac{1}{20})$ $\epsilon\times(2,4,8)$ $N\times(10,10^{2},10^{3})$
1.51,1.63, ${\bf 1.51},1.63,\,1.65$ 1.40, 1.14, 1.071.401.141.071.40,\,1.14,\,1.071.40 , 1.14 , 1.07 1.27, 1.10, 1.031.271.101.031.27,\,1.10,\,1.031.27 , 1.10 , 1.03 $1.14,1.02,1.00$
Table 4: Ratio between $\hat{\sigma}_{\rm MC}$ and $\hat{\sigma}_{\rm H}/\sqrt{k}$ for the toy model of section 2 by varying one configuration parameter at the time. The value for the nominal model is given by the number in bold.
3.7 Adding external constraints
In section 3.1, all parameters of the likelihood were assumed to be unconstrained, though this is certainly not the most general case of interest. Indeed, it often happens that at least some of the nuisance parameters can be constrained by independent auxiliary measurements. Let us then assume that the likelihood function includes both constrained and unconstrained nuisance parameters. If the set of $q$ constrained nuisance parameters is denoted by $\boldsymbol{\tau}$, the right-hand side of eq. (25) gets modified to
${\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau})={\bf f}_{0}+{\bf j}(\mu-\mu_{0})+{\bf J}(\boldsymbol{\theta}-\boldsymbol{\theta}_{0})+{\bf K}(\boldsymbol{\tau}-\boldsymbol{\tau}_{0})+\cdots,$ (62)
where ${\bf K}$ is the $n\times q$ Jacobian matrix associated with the NP subject to a constraint. In some cases, the effect of a constrained nuisance parameter relative to the nominal prediction is known exactly: for example, a luminosity uncertainty on the total MC normalization can be modeled as a multiplicative correction, i.e. ${\bf K}\propto{\bf f}_{0}$. In other cases, however, ${\bf K}$ needs to be estimated numerically, giving rise to an additional source of systematic uncertainty, if MC samples of finite size are used to determine its elements. The negative log-likelihood function in eq. (24) should be then extended accordingly to
$-\ln{L}(\mu,\boldsymbol{\theta},\boldsymbol{\tau})$ $\approx\frac{1}{2}\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau})\right)^{T}{\bf V}^{-1}\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau})\right)$
$+\frac{1}{2}(\boldsymbol{\tau}-\boldsymbol{\tau}_{0})^{T}(\boldsymbol{\tau}-\boldsymbol{\tau}_{0})+{\rm const.}$ (63)
where the estimators from the auxiliary measurements have been assumed as normally-distributed and independent from each other Pinto:2023yob . The additional parameters $\boldsymbol{\tau}$ can be profiled at a fixed value of $(\mu,\boldsymbol{\theta})$ yielding the well-known result Pinto:2023yob ; Behnke :
$-2\ln\frac{L(\mu,\boldsymbol{\theta},\hat{\boldsymbol{\tau}}_{(\mu,\boldsymbol{\theta}}))}{{L}(\hat{\mu},\hat{\boldsymbol{\theta}},\hat{\boldsymbol{\tau}})}=\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau}_{0})\right)^{T}\left({\bf V}+{\bf K}{\bf K}^{T}\right)^{-1}\left({\bf y}-{\bf f}(\mu,\boldsymbol{\theta},\boldsymbol{\tau}_{0})\right),$ (64)
that is, the Poisson covariance matrix of the data ${\bf V}$ is augmented by the non-diagonal matrix ${\bf K}{\bf K}^{T}$, with their sum still symmetric and positive-definite. By means of Woodbury’s matrix identity, we can expand
$\left({\bf V}+{\bf K}{\bf K}^{T}\right)^{-1}$ $={\bf V}^{-\frac{1}{2}}\left({\bf 1}-{\bf H}\left(1+{\bf H}^{T}{\bf H}\right)^{-1}{\bf H}^{T}\right){\bf V}^{-\frac{1}{2}}$
$={\bf V}^{-\frac{1}{2}}{\bf Z}^{-1}{\bf V}^{-\frac{1}{2}},$ (65)
with ${\bf H}={\bf V}^{-\frac{1}{2}}{\bf K}$. The newly introduced matrix ${\bf Z}$ is a symmetric and positive-definite matrix, so it is invertible and has a unique square-root matrix ${\bf Z}^{\frac{1}{2}}$. By means of eq. (3.7), eq. (64) can be ultimately recast in the same form of eq. (26), namely
$\lVert{\bf r}_{\mu}^{\prime}-{\bf A}^{\prime}(\boldsymbol{\theta}-\boldsymbol{\theta}_{0})\rVert^{2},$ (66)
where ${\bf r}_{\mu}^{\prime}$ and ${\bf A}^{\prime}$ are defined as in eq. (27) with the replacement ${\bf V}^{-1/2}\to\left({\bf V}{\bf Z}\right)^{-1/2}$. In this framework, statistical fluctuations affecting the entries of the Jacobian matrix ${\bf K}$ are propagated into the NLL function by the matrix ${\bf Z}$ which mixes them with fluctuations of ${\bf J}$ in a possibly intricate way. However, due to the formal identity between eq. (26) and (66), the same analysis pursued in section 3.2 still holds. Notice, however, that in the case of eq. (3.7) some of the scaling laws of eq. (55) are broken due to a non-trivial combination of statistical and systematic uncertainties. Also, the inclusion of external constraints in the likelihood function is likely to reduce the global correlation factor of the POI and hence improve the validity of the asymptotic approximation.
Constrained nuisance parameters naturally arise in problems where the nominal model prediction is affected by sources of systematic uncertainty for which some a priori knowledge is available from independent measurements or theoretical prejudice. In this respect, $K_{ij}$ can be seen as the “error” on the expectation value in the $i^{\rm th}$ bin as due to a ±$\pm 1\sigma$ “error” of the $j^{\rm th}$ source of systematic uncertainty. Thus, when ${\bf K}$ is modeled via MC simulation, statistical fluctuations of the MC sample induce “errors on the errors”. This subject has a connection with ref. Cowan:2018lhq , where breaking of asymptotic properties of the MLE has been also observed.
4 Conclusions
In this work, we have studied the problem of setting confidence intervals on a parameter of interest measured from the binned analysis of high-statistics and high-granularity data sets, focusing on the case where the likelihood function depends on additional nuisance parameters and is constructed from finite-size MC samples. We have set ourselves in the limit of large event counts in all bins for both the data and the MC-simulated samples, such that asymptotic formulas from the theory of maximum-likelihood estimators would be naively expected to hold. In particular, by assuming Wilks’ theorem, central confidence intervals can be defined from either the profile-likelihood ratio or the Hessian matrix.
We have designed a simple toy model consisting of a signal and background process where the unknown signal (background) normalization is treated as the parameter of interest (nuisance parameter). We have further assumed that the likelihood function is constructed using Monte Carlo simulations of finite size. This toy model depends on configuration parameters that we are free to choose: the total number of expected events, the number of bins, a model parameter that increases the signal-to-background separation, and the statistical power of the MC samples relative to the data. By inspecting a large number of pseudo-experiments, we could establish numerically that asymptotic formulas are broken even when the statistical power of the simulated samples is a few times larger than that of the data. In particular, confidence intervals based on Wilks’ theorem suffer from systematic under-coverage. Consequently, we have tried alternative methods to set confidence intervals based on the post-fit distribution of the profile-likelihood ratio, such as the profiled Feldman-Cousins approach or variants thereof acero2024 ; Barlett ; COUSINS1992331 , but none of them succeeded in recovering the correct coverage.
Methods based on the asymptotic properties of the maximum-likelihood estimator ultimately yield the right coverage, but only in the limit of a very large scaling of the data and/or simulated sample size, though we could not find a first-principle indication of how large such a scaling should be. Finally, by combining the Hessian uncertainties obtained from different likelihood functions, we have worked out a heuristic prescription for a confidence interval, defined in eq. (22), which gets closer to the correct Feldman-Cousins interval for this problem, thus providing a practical test to quantify the level of under-coverage when other standard confidence intervals are used.
Inspired by this toy model, we have studied the problem from a more general perspective, with the goal of understanding in which circumstances statistical fluctuations in the MC prediction can reduce the uncertainty predicted for the parameter of interest, ultimately leading to under-coverage of confidence intervals constructed in the asymptotic approximation. We found that this artificially augmented sensitivity can be expressed in terms of a bias term to the quadratic form $S$ defined in eq. (32). We have identified and characterized two main sources of a positive bias, which then leads to under-coverage: fluctuations on the gradient of the model function with respect to the POI, and fluctuations on the Jacobian matrix with respect to the nuisance parameters. In the latter case, the effect can be greatly enhanced in the presence of a large global correlation of the parameter of interest, as defined by eq. (38). Both effects can explain the results of the toy study. Interestingly, the bias term is not proportional to $S$ by a naive factor of $k^{-1}$, suggesting that extensions of the likelihood functions that include the effect of finite MC samples in an effective way, such as the popular Barlow-Beeston lite method BBlite , can be insufficient to provide a full account of these effects. Based on general scaling laws, we propose a practical test to estimate the standard deviation that the MLE for the parameter of interest would have in the limit of infinite MC statistics. As the latter should be smaller compared to the actual statistical spread when finite-MC samples are used, it could be used as a reference to gauge how far an analysis is from the asymptotic regime.
While Wilks’ theorem for the profile-likelihood ratio guarantees that the desired asymptotic behavior will eventually be attained, the assumption of being in the asymptotic regime could be wrong even when the event counts are large in absolute terms for both data and simulation: this suggests that the MC statistical uncertainty that needs to be “small” to be safely in the asymptotic regime is not necessarily provided by total number of observed/simulated events bin, nor by the total number of bins. Care should be therefore taken when reporting the uncertainty on parameters of interest based on asymptotic formulas when the data model is derived from finite-size MC samples and additional nuisance parameters are profiled. This seems to be the case also for high-statistics experiments, and even when the statistical power of the simulated samples is comparable to that of the data.
Appendix A Approximate likelihood functions for the toy model
The likelihood functions described below encode different levels of approximating eq. (7). For each definition, it should be understood that $\mu_{1}$ and $\mu_{2}$ are the appropriate functions $\mu$ and $\theta$, depending on whether the original or transformed parameters are being considered.
  • Poisson without MC stat.
    $-\ln L_{\rm P}(\mu,\theta)=\sum_{i=1}^{n}\left[\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)-y_{i}\ln\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)\right]$ (67)
    This would be the correct likelihood function if the MC predictions $t_{ji}$ were exact. For finite MC samples, it must be considered as a crude approximation.
  • Gauss without MC stat.
    $-\ln L_{\rm G}(\mu,\theta)=\frac{1}{2}\sum_{i=1}^{n}\frac{\left(y_{i}-\left(\mu_{1}\frac{t_{1k}}{k_{1}}+\mu_{2}\frac{t_{2k}}{k_{2}}\right)\right)^{2}}{y_{i}}.$ (68)
    This function provides an excellent approximation of the Poisson likelihood in eq. (67) in the limit $y_{i}\gg 1$. An interesting feature of eq. (68) is that its minimum can be found analytically by solving a linear system of equations Cowan .
  • Gauss + MC stat.
    $-\ln L_{\rm G+MC}(\mu,\theta)=\frac{1}{2}\sum_{i=1}^{n}\frac{\left(y_{i}-\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{1}}\right)\right)^{2}}{y_{i}+\frac{t_{1i}}{k_{1}^{2}}+\frac{t_{2i}}{k_{2}^{2}}}.$ (69)
    This is the simplest extension of eq. (68) that accounts for statistical uncertainties on the MC templates: the total MC variance per bin is added to the Poisson variance of the data. The global minimum can still be found analytically.
  • Gauss + Barlow-Beeston lite
    $-\ln L_{\rm G+BBlite}(\mu,\theta,\boldsymbol{\beta})=$ (70)
    $\frac{1}{2}\sum_{i=1}^{n}\left[\frac{\left(y_{i}-\beta_{i}\cdot\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)\right)^{2}}{y_{i}}+m_{i}\cdot\left(\beta_{i}-1\right)^{2}\right].$
    This is a well-known and widely used approximation of the Barlow-Beeston likelihood BBlite : it consists in assigning a unique parameter $\beta_{i}$ for all MC processes in each bin constrained by the total effective relative uncertainty $1/\sqrt{m_{i}}$.
  • Gauss + Barlow-Beeston
    $-\ln L_{\rm G+BB}(\mu,\theta,\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=$ (71)
    $\frac{1}{2}\sum_{i=1}^{n}\left[\frac{\left(y_{i}-\left(\mu_{1}\frac{\theta_{1i}}{k_{1}}+\mu_{2}\frac{\theta_{2i}}{k_{2}}\right)\right)^{2}}{y_{i}}+\sum_{j=1}^{2}\frac{\left(\theta_{ji}-t_{ji}\right)^{2}}{t_{ji}}\right].$
    This is the Barlow-Beeston likelihood of eq. (7) in the Gaussian approximation.
In the likelihood functions defined by eq. (70) and (71), the nuisance parameters related to the MC templates can be profiled analytically. For example, for eq. (70) it is easy to prove that:
$(\hat{\mu},\hat{\theta})=\underset{\mu,\theta}{\rm argmin}\left[\sum_{i=1}^{n}\frac{\left(y_{i}-\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)\right)^{2}}{y_{i}+\frac{\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)^{2}}{m_{i}}}\right]$ (72)
while for eq. (71) we get:
$(\hat{\mu},\hat{\theta})=\underset{\mu,\theta}{\rm argmin}\left[\sum_{i=1}^{n}\frac{\left(y_{i}-\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)\right)^{2}}{y_{i}+\mu_{1}^{2}\frac{t_{1i}}{k_{1}^{2}}+\mu_{2}^{2}\frac{t_{2i}}{k_{2}^{2}}}\right]$ (73)
It has been verified that minimizing eq. (71) over the full set of parameters or solving eq. (73) leads to identical results within numerical accuracy.
Appendix B Numerical results for alternative model configurations
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
Poisson Numeric Hessian 0 ${\bf 0.661}(5)$ 0.087, $,\;0.087$ 0.087
$+5\bar{\sigma}_{{\rm H}}$ $0.529(5)$ 0.087, $,\;0.087$ 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.087, $,\;0.087$ 0.087
Gauss Numeric Hessian 0 ${\bf 0.663}(5)$ 0.087, $,\;0.087$ 0.087
$+5\bar{\sigma}_{{\rm H}}$ $0.528(5)$ 0.087, $,\;0.087$ 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.087, $,\;0.087$ 0.087
Gauss Analytic Hessian 0 ${\bf 0.663}(5)$ 0.087, $,\;0.087$ 0.087
$+5\bar{\sigma}_{{\rm H}}$ $0.528(5)$ 0.087, $,\;0.087$ 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.087, $,\;0.087$ 0.087
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.683}(5)$ 0.091, $,\;0.091$ 0.091
$+5\bar{\sigma}_{{\rm H}}$ $0.552(5)$ 0.091, $,\;0.091$ 0.091
$-5\bar{\sigma}_{{\rm H}}$ $0.554(5)$ 0.091, $,\;0.091$ 0.091
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.683}(5)$ 0.091, $,\;0.091$ 0.091
$+5\bar{\sigma}_{{\rm H}}$ $0.552(5)$ 0.091, $,\;0.091$ 0.091
$-5\bar{\sigma}_{{\rm H}}$ $0.554(5)$ 0.091, $,\;0.091$ 0.091
Gauss + BB Numeric Hessian 0 $0.647(5)$ 0.099, $,\;0.099$ 0.099
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.649}(5)$ 0.100, $,\;0.100$ 0.100
$-5\bar{\sigma}_{{\rm H}}$ $0.638(5)$ 0.100, $,\;0.100$ 0.100
Gauss + BB Numeric PLR 0 $0.647(5)$ 0.099, $,\;0.099$ 0.099
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.650}(5)$ 0.100, $,\;0.100$ 0.100
$-5\bar{\sigma}_{{\rm H}}$ $0.639(5)$ 0.100, $,\;0.100$ 0.100
Gauss + BB Numeric sPFC 0 ${\bf 0.662}(5)$ 0.103, $,\;0.103$ 0.103
$+5\bar{\sigma}_{{\rm H}}$ $0.745(4)$ 0.124, 0.1230.1240.1230.124,\;0.1230.124 , 0.123
$-5\bar{\sigma}_{{\rm H}}$ $0.746(4)$ 0.123, 0.1220.1230.1220.123,\;0.1220.123 , 0.122
Gauss + BB Numeric PFC 0 ${\bf 0.654}(5)$ 0.102, $,\;0.102$ 0.102
$+5\bar{\sigma}_{{\rm H}}$ $0.753(4)$ 0.124, 0.1230.1240.1230.124,\;0.1230.124 , 0.123
$-5\bar{\sigma}_{{\rm H}}$ $0.747(4)$ 0.123, 0.1220.1230.1220.123,\;0.1220.123 , 0.122
Gauss + BB Numeric Barlett 0 ${\bf 0.653}(5)$ 0.102, $,\;0.102$ 0.102
$+5\bar{\sigma}_{{\rm H}}$ $0.740(4)$ 0.121, $,\;0.121$ 0.121
$-5\bar{\sigma}_{{\rm H}}$ $0.738(4)$ 0.121, 0.1200.1210.1200.121,\;0.1200.121 , 0.120
Gauss + BB Numeric CH 0 ${\bf 0.679}(5)$ 0.107, $,\;0.107$ 0.107
$+5\bar{\sigma}_{{\rm H}}$ $0.788(4)$ 0.135, 0.1340.1350.1340.135,\;0.1340.135 , 0.134
$-5\bar{\sigma}_{{\rm H}}$ $0.785(4)$ 0.134, 0.1330.1340.1330.134,\;0.1330.134 , 0.133
Gauss + BB Numeric FC Cheat 0 ${\bf 0.680}(5)$ 0.108, 0.1060.1080.1060.108,\;0.1060.108 , 0.106
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.109, 0.1070.1090.1070.109,\;0.1070.109 , 0.107
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.108, 0.1060.1080.1060.108,\;0.1060.108 , 0.106
Gauss + BB Numeric Heuristic 0 ${\bf 0.683}(5)$ 0.107, $,\;0.107$ 0.107
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.686}(5)$ 0.108, $,\;0.108$ 0.108
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.677}(5)$ 0.109, $,\;0.109$ 0.109
Table 5: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=10$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
Poisson Numeric Hessian 0 ${\bf 0.669}(5)$ 0.092, $,\;0.092$ 0.092
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.092, $,\;0.092$ 0.092
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.665}(5)$ 0.092, $,\;0.092$ 0.092
Gauss Numeric Hessian 0 ${\bf 0.669}(5)$ 0.092, $,\;0.092$ 0.092
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.092, $,\;0.092$ 0.092
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.667}(5)$ 0.092, $,\;0.092$ 0.092
Gauss Analytic Hessian 0 ${\bf 0.669}(5)$ 0.092, $,\;0.092$ 0.092
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.092, $,\;0.092$ 0.092
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.667}(5)$ 0.092, $,\;0.092$ 0.092
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.675}(5)$ 0.093, $,\;0.093$ 0.093
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.679}(5)$ 0.093, $,\;0.093$ 0.093
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.094, $,\;0.094$ 0.094
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.675}(5)$ 0.093, $,\;0.093$ 0.093
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.678}(5)$ 0.093, $,\;0.093$ 0.093
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.094, $,\;0.094$ 0.094
Gauss + BB Numeric Hessian 0 ${\bf 0.665}(5)$ 0.095, $,\;0.095$ 0.095
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.674}(5)$ 0.096, $,\;0.096$ 0.096
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.668}(5)$ 0.096, $,\;0.096$ 0.096
Gauss + BB Numeric PLR 0 ${\bf 0.665}(5)$ 0.095, $,\;0.095$ 0.095
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.675}(5)$ 0.096, $,\;0.096$ 0.096
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.669}(5)$ 0.096, $,\;0.096$ 0.096
Gauss + BB Numeric sPFC 0 ${\bf 0.679}(5)$ 0.096, $,\;0.096$ 0.096
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.689}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.679}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric PFC 0 ${\bf 0.675}(5)$ 0.096, $,\;0.096$ 0.096
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.682}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric Barlett 0 ${\bf 0.675}(5)$ 0.096, $,\;0.096$ 0.096
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.681}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric CH 0 ${\bf 0.681}(5)$ 0.098, $,\;0.098$ 0.098
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.688}(5)$ 0.100, $,\;0.100$ 0.100
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.698}(5)$ 0.100, $,\;0.100$ 0.100
Gauss + BB Numeric FC Cheat 0 ${\bf 0.683}(5)$ 0.098, 0.0970.0980.0970.098,\;0.0970.098 , 0.097
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.689}(5)$ 0.098, 0.0970.0980.0970.098,\;0.0970.098 , 0.097
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric Heuristic 0 ${\bf 0.675}(5)$ 0.098, $,\;0.098$ 0.098
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.683}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.678}(5)$ 0.098, $,\;0.098$ 0.098
Table 6: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=40$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.047, $,\;0.047$ 0.047
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.047, $,\;0.047$ 0.047
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.047, $,\;0.047$ 0.047
Poisson Numeric Hessian 0 $0.519(5)$ 0.039, $,\;0.039$ 0.039
$+5\bar{\sigma}_{{\rm H}}$ $0.248(4)$ 0.039, $,\;0.039$ 0.039
$-5\bar{\sigma}_{{\rm H}}$ $0.254(4)$ 0.039, $,\;0.039$ 0.039
Gauss Numeric Hessian 0 $0.520(5)$ 0.039, $,\;0.039$ 0.039
$+5\bar{\sigma}_{{\rm H}}$ $0.249(4)$ 0.039, $,\;0.039$ 0.039
$-5\bar{\sigma}_{{\rm H}}$ $0.255(4)$ 0.039, $,\;0.039$ 0.039
Gauss Analytic Hessian 0 $0.520(5)$ 0.039, $,\;0.039$ 0.039
$+5\bar{\sigma}_{{\rm H}}$ $0.249(4)$ 0.039, $,\;0.039$ 0.039
$-5\bar{\sigma}_{{\rm H}}$ $0.255(4)$ 0.039, $,\;0.039$ 0.039
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.680}(5)$ 0.056, $,\;0.056$ 0.056
$+5\bar{\sigma}_{{\rm H}}$ $0.362(5)$ 0.056, $,\;0.056$ 0.056
$-5\bar{\sigma}_{{\rm H}}$ $0.372(5)$ 0.056, $,\;0.056$ 0.056
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.680}(5)$ 0.056, $,\;0.056$ 0.056
$+5\bar{\sigma}_{{\rm H}}$ $0.363(5)$ 0.055, $,\;0.055$ 0.055
$-5\bar{\sigma}_{{\rm H}}$ $0.372(5)$ 0.056, $,\;0.056$ 0.056
Gauss + BB Numeric Hessian 0 $0.592(5)$ 0.067, $,\;0.067$ 0.067
$+5\bar{\sigma}_{{\rm H}}$ $0.600(5)$ 0.068, 0.0670.0680.0670.068,\;0.0670.068 , 0.067
$-5\bar{\sigma}_{{\rm H}}$ $0.598(5)$ 0.068, $,\;0.068$ 0.068
Gauss + BB Numeric PLR 0 $0.593(5)$ 0.067, $,\;0.067$ 0.067
$+5\bar{\sigma}_{{\rm H}}$ $0.602(5)$ 0.068, $,\;0.068$ 0.068
$-5\bar{\sigma}_{{\rm H}}$ $0.599(5)$ 0.068, 0.0670.0680.0670.068,\;0.0670.068 , 0.067
Gauss + BB Numeric sPFC 0 $0.500(5)$ 0.055, 0.0540.0550.0540.055,\;0.0540.055 , 0.054
$+5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.060, 0.0590.0600.0590.060,\;0.0590.060 , 0.059
$-5\bar{\sigma}_{{\rm H}}$ $0.528(5)$ 0.059, $,\;0.059$ 0.059
Gauss + BB Numeric PFC 0 $0.560(5)$ 0.067, 0.0610.0670.0610.067,\;0.0610.067 , 0.061
$+5\bar{\sigma}_{{\rm H}}$ $0.628(5)$ 0.072, 0.0640.0720.0640.072,\;0.0640.072 , 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.628(5)$ 0.071, 0.0630.0710.0630.071,\;0.0630.071 , 0.063
Gauss + BB Numeric Barlett 0 $0.553(5)$ 0.065, 0.0610.0650.0610.065,\;0.0610.065 , 0.061
$+5\bar{\sigma}_{{\rm H}}$ $0.598(5)$ 0.070, 0.0630.0700.0630.070,\;0.0630.070 , 0.063
$-5\bar{\sigma}_{{\rm H}}$ $0.595(5)$ 0.069, 0.0620.0690.0620.069,\;0.0620.069 , 0.062
Gauss + BB Numeric CH 0 ${\bf 0.687}(5)$ 0.081, 0.0770.0810.0770.081,\;0.0770.081 , 0.077
$+5\bar{\sigma}_{{\rm H}}$ $0.794(4)$ 0.099, 0.0900.0990.0900.099,\;0.0900.099 , 0.090
$-5\bar{\sigma}_{{\rm H}}$ $0.789(4)$ 0.097, 0.0880.0970.0880.097,\;0.0880.097 , 0.088
Gauss + BB Numeric FC Cheat 0 ${\bf 0.691}(5)$ 0.081, 0.0710.0810.0710.081,\;0.0710.081 , 0.071
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.681}(5)$ 0.083, 0.0720.0830.0720.083,\;0.0720.083 , 0.072
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.678}(5)$ 0.081, 0.0700.0810.0700.081,\;0.0700.081 , 0.070
Gauss + BB Numeric Heuristic 0 ${\bf 0.682}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.081, $,\;0.081$ 0.081
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.082, 0.0810.0820.0810.082,\;0.0810.082 , 0.081
Table 7: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.06$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.024, $,\;0.024$ 0.024
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.024, $,\;0.024$ 0.024
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.024, $,\;0.024$ 0.024
Poisson Numeric Hessian 0 $0.523(5)$ 0.022, $,\;0.022$ 0.022
$+5\bar{\sigma}_{{\rm H}}$ $0.485(5)$ 0.022, $,\;0.022$ 0.022
$-5\bar{\sigma}_{{\rm H}}$ $0.498(5)$ 0.022, $,\;0.022$ 0.022
Gauss Numeric Hessian 0 $0.522(5)$ 0.022, $,\;0.022$ 0.022
$+5\bar{\sigma}_{{\rm H}}$ $0.486(5)$ 0.022, $,\;0.022$ 0.022
$-5\bar{\sigma}_{{\rm H}}$ $0.496(5)$ 0.022, $,\;0.022$ 0.022
Gauss Analytic Hessian 0 $0.522(5)$ 0.022, $,\;0.022$ 0.022
$+5\bar{\sigma}_{{\rm H}}$ $0.486(5)$ 0.022, $,\;0.022$ 0.022
$-5\bar{\sigma}_{{\rm H}}$ $0.496(5)$ 0.022, $,\;0.022$ 0.022
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.689}(5)$ 0.032, $,\;0.032$ 0.032
$+5\bar{\sigma}_{{\rm H}}$ $0.648(5)$ 0.032, $,\;0.032$ 0.032
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.655}(5)$ 0.032, $,\;0.032$ 0.032
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.688}(5)$ 0.032, $,\;0.032$ 0.032
$+5\bar{\sigma}_{{\rm H}}$ $0.648(5)$ 0.032, $,\;0.032$ 0.032
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.656}(5)$ 0.032, $,\;0.032$ 0.032
Gauss + BB Numeric Hessian 0 ${\bf 0.663}(5)$ 0.033, $,\;0.033$ 0.033
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.658}(5)$ 0.033, $,\;0.033$ 0.033
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.661}(5)$ 0.034, $,\;0.034$ 0.034
Gauss + BB Numeric PLR 0 ${\bf 0.664}(5)$ 0.033, $,\;0.033$ 0.033
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.658}(5)$ 0.034, 0.0330.0340.0330.034,\;0.0330.034 , 0.033
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.662}(5)$ 0.034, $,\;0.034$ 0.034
Gauss + BB Numeric sPFC 0 $0.521(5)$ 0.025, $,\;0.025$ 0.025
$+5\bar{\sigma}_{{\rm H}}$ $0.512(5)$ 0.025, $,\;0.025$ 0.025
$-5\bar{\sigma}_{{\rm H}}$ $0.519(5)$ 0.025, $,\;0.025$ 0.025
Gauss + BB Numeric PFC 0 $0.596(5)$ 0.030, 0.0270.0300.0270.030,\;0.0270.030 , 0.027
$+5\bar{\sigma}_{{\rm H}}$ $0.593(5)$ 0.030, 0.0280.0300.0280.030,\;0.0280.030 , 0.028
$-5\bar{\sigma}_{{\rm H}}$ $0.592(5)$ 0.031, 0.0280.0310.0280.031,\;0.0280.031 , 0.028
Gauss + BB Numeric Barlett 0 $0.588(5)$ 0.029, 0.0270.0290.0270.029,\;0.0270.029 , 0.027
$+5\bar{\sigma}_{{\rm H}}$ $0.580(5)$ 0.030, 0.0270.0300.0270.030,\;0.0270.030 , 0.027
$-5\bar{\sigma}_{{\rm H}}$ $0.579(5)$ 0.030, 0.0270.0300.0270.030,\;0.0270.030 , 0.027
Gauss + BB Numeric CH 0 ${\bf 0.700}(5)$ 0.035, 0.0330.0350.0330.035,\;0.0330.035 , 0.033
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.711}(5)$ 0.036, 0.0330.0360.0330.036,\;0.0330.036 , 0.033
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.706}(5)$ 0.036, 0.0330.0360.0330.036,\;0.0330.036 , 0.033
Gauss + BB Numeric FC Cheat 0 ${\bf 0.683}(5)$ 0.035, 0.0310.0350.0310.035,\;0.0310.035 , 0.031
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.683}(5)$ 0.036, 0.0310.0360.0310.036,\;0.0310.036 , 0.031
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.036, 0.0310.0360.0310.036,\;0.0310.036 , 0.031
Gauss + BB Numeric Heuristic 0 ${\bf 0.689}(5)$ 0.035, $,\;0.035$ 0.035
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.685}(5)$ 0.035, $,\;0.035$ 0.035
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.689}(5)$ 0.035, $,\;0.035$ 0.035
Table 8: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.12$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
Poisson Numeric Hessian 0 $0.530(5)$ 0.069, $,\;0.069$ 0.069
$+5\bar{\sigma}_{{\rm H}}$ $0.060(2)$ 0.069, $,\;0.069$ 0.069
$-5\bar{\sigma}_{{\rm H}}$ $0.067(2)$ 0.069, $,\;0.069$ 0.069
Gauss Numeric Hessian 0 $0.531(5)$ 0.069, $,\;0.069$ 0.069
$+5\bar{\sigma}_{{\rm H}}$ $0.060(2)$ 0.069, $,\;0.069$ 0.069
$-5\bar{\sigma}_{{\rm H}}$ $0.067(3)$ 0.069, $,\;0.069$ 0.069
Gauss Analytic Hessian 0 $0.531(5)$ 0.069, $,\;0.069$ 0.069
$+5\bar{\sigma}_{{\rm H}}$ $0.060(2)$ 0.069, $,\;0.069$ 0.069
$-5\bar{\sigma}_{{\rm H}}$ $0.067(3)$ 0.069, $,\;0.069$ 0.069
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.691}(5)$ 0.098, 0.0970.0980.0970.098,\;0.0970.098 , 0.097
$+5\bar{\sigma}_{{\rm H}}$ $0.107(3)$ 0.098, 0.0970.0980.0970.098,\;0.0970.098 , 0.097
$-5\bar{\sigma}_{{\rm H}}$ $0.110(3)$ 0.098, $,\;0.098$ 0.098
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.692}(5)$ 0.098, 0.0970.0980.0970.098,\;0.0970.098 , 0.097
$+5\bar{\sigma}_{{\rm H}}$ $0.107(3)$ 0.097, $,\;0.097$ 0.097
$-5\bar{\sigma}_{{\rm H}}$ $0.110(3)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric Hessian 0 $0.545(5)$ 0.137, 0.1340.1370.1340.137,\;0.1340.137 , 0.134
$+5\bar{\sigma}_{{\rm H}}$ $0.549(5)$ 0.144, 0.1400.1440.1400.144,\;0.1400.144 , 0.140
$-5\bar{\sigma}_{{\rm H}}$ $0.548(5)$ 0.145, 0.1410.1450.1410.145,\;0.1410.145 , 0.141
Gauss + BB Numeric PLR 0 $0.552(5)$ 0.138, 0.1340.1380.1340.138,\;0.1340.138 , 0.134
$+5\bar{\sigma}_{{\rm H}}$ $0.557(5)$ 0.150, 0.1450.1500.1450.150,\;0.1450.150 , 0.145
$-5\bar{\sigma}_{{\rm H}}$ $0.556(5)$ 0.141, 0.1370.1410.1370.141,\;0.1370.141 , 0.137
Gauss + BB Numeric sPFC 0 $0.481(5)$ 0.126, 0.1210.1260.1210.126,\;0.1210.126 , 0.121
$+5\bar{\sigma}_{{\rm H}}$ $0.537(5)$ 0.149, 0.1430.1490.1430.149,\;0.1430.149 , 0.143
$-5\bar{\sigma}_{{\rm H}}$ $0.543(5)$ 0.138, 0.1340.1380.1340.138,\;0.1340.138 , 0.134
Gauss + BB Numeric PFC 0 $0.529(5)$ 0.151, 0.1380.1510.1380.151,\;0.1380.151 , 0.138
$+5\bar{\sigma}_{{\rm H}}$ $0.637(5)$ 0.181, 0.1530.1810.1530.181,\;0.1530.181 , 0.153
$-5\bar{\sigma}_{{\rm H}}$ $0.635(5)$ 0.169, 0.1450.1690.1450.169,\;0.1450.169 , 0.145
Gauss + BB Numeric Barlett 0 $0.523(5)$ 0.147, 0.1360.1470.1360.147,\;0.1360.147 , 0.136
$+5\bar{\sigma}_{{\rm H}}$ $0.599(5)$ 0.172, 0.1490.1720.1490.172,\;0.1490.172 , 0.149
$-5\bar{\sigma}_{{\rm H}}$ $0.594(5)$ 0.161, 0.1410.1610.1410.161,\;0.1410.161 , 0.141
Gauss + BB Numeric CH 0 ${\bf 0.692}(5)$ 0.193, 0.1800.1930.1800.193,\;0.1800.193 , 0.180
$+5\bar{\sigma}_{{\rm H}}$ $0.815(4)$ 0.266, 0.2490.2660.2490.266,\;0.2490.266 , 0.249
$-5\bar{\sigma}_{{\rm H}}$ $0.821(4)$ 0.253, 0.2370.2530.2370.253,\;0.2370.253 , 0.237
Gauss + BB Numeric FC Cheat 0 ${\bf 0.688}(5)$ 0.191, 0.1700.1910.1700.191,\;0.1700.191 , 0.170
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.683}(5)$ 0.199, 0.1740.1990.1740.199,\;0.1740.199 , 0.174
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.683}(5)$ 0.186, 0.1640.1860.1640.186,\;0.1640.186 , 0.164
Gauss + BB Numeric Heuristic 0 ${\bf 0.691}(5)$ 0.186, 0.1830.1860.1830.186,\;0.1830.186 , 0.183
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.697}(5)$ 0.196, 0.1920.1960.1920.196,\;0.1920.196 , 0.192
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.695}(5)$ 0.197, 0.1920.1970.1920.197,\;0.1920.197 , 0.192
Table 9: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=100$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
Poisson Numeric Hessian 0 $0.514(5)$ 0.088, $,\;0.088$ 0.088
$+5\bar{\sigma}_{{\rm H}}$ $0.450(5)$ 0.088, 0.0870.0880.0870.088,\;0.0870.088 , 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.442(5)$ 0.088, $,\;0.088$ 0.088
Gauss Numeric Hessian 0 $0.514(5)$ 0.088, $,\;0.088$ 0.088
$+5\bar{\sigma}_{{\rm H}}$ $0.450(5)$ 0.088, 0.0870.0880.0870.088,\;0.0870.088 , 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.441(5)$ 0.088, $,\;0.088$ 0.088
Gauss Analytic Hessian 0 $0.514(5)$ 0.088, $,\;0.088$ 0.088
$+5\bar{\sigma}_{{\rm H}}$ $0.450(5)$ 0.088, 0.0870.0880.0870.088,\;0.0870.088 , 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.441(5)$ 0.088, $,\;0.088$ 0.088
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.676}(5)$ 0.125, 0.1240.1250.1240.125,\;0.1240.125 , 0.124
$+5\bar{\sigma}_{{\rm H}}$ $0.604(5)$ 0.125, 0.1240.1250.1240.125,\;0.1240.125 , 0.124
$-5\bar{\sigma}_{{\rm H}}$ $0.593(5)$ 0.125, 0.1240.1250.1240.125,\;0.1240.125 , 0.124
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.676}(5)$ 0.125, 0.1240.1250.1240.125,\;0.1240.125 , 0.124
$+5\bar{\sigma}_{{\rm H}}$ $0.603(5)$ 0.124, $,\;0.124$ 0.124
$-5\bar{\sigma}_{{\rm H}}$ $0.594(5)$ 0.125, 0.1240.1250.1240.125,\;0.1240.125 , 0.124
Gauss + BB Numeric Hessian 0 $0.641(5)$ 0.136, 0.1340.1360.1340.136,\;0.1340.136 , 0.134
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.653}(5)$ 0.143, 0.1410.1430.1410.143,\;0.1410.143 , 0.141
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.653}(5)$ 0.144, 0.1410.1440.1410.144,\;0.1410.144 , 0.141
Gauss + BB Numeric PLR 0 $0.649(5)$ 0.136, 0.1350.1360.1350.136,\;0.1350.136 , 0.135
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.660}(5)$ 0.148, 0.1450.1480.1450.148,\;0.1450.148 , 0.145
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.660}(5)$ 0.140, 0.1380.1400.1380.140,\;0.1380.140 , 0.138
Gauss + BB Numeric sPFC 0 $0.511(5)$ 0.101, 0.1000.1010.1000.101,\;0.1000.101 , 0.100
$+5\bar{\sigma}_{{\rm H}}$ $0.501(5)$ 0.104, 0.1020.1040.1020.104,\;0.1020.104 , 0.102
$-5\bar{\sigma}_{{\rm H}}$ $0.495(5)$ 0.100, 0.0990.1000.0990.100,\;0.0990.100 , 0.099
Gauss + BB Numeric PFC 0 $0.569(5)$ 0.127, 0.1160.1270.1160.127,\;0.1160.127 , 0.116
$+5\bar{\sigma}_{{\rm H}}$ $0.588(5)$ 0.138, 0.1200.1380.1200.138,\;0.1200.138 , 0.120
$-5\bar{\sigma}_{{\rm H}}$ $0.592(5)$ 0.130, 0.1150.1300.1150.130,\;0.1150.130 , 0.115
Gauss + BB Numeric Barlett 0 $0.559(5)$ 0.123, 0.1150.1230.1150.123,\;0.1150.123 , 0.115
$+5\bar{\sigma}_{{\rm H}}$ $0.568(5)$ 0.133, 0.1180.1330.1180.133,\;0.1180.133 , 0.118
$-5\bar{\sigma}_{{\rm H}}$ $0.572(5)$ 0.125, 0.1140.1250.1140.125,\;0.1140.125 , 0.114
Gauss + BB Numeric CH 0 ${\bf 0.683}(5)$ 0.149, 0.1400.1490.1400.149,\;0.1400.149 , 0.140
$+5\bar{\sigma}_{{\rm H}}$ $0.724(4)$ 0.166, 0.1460.1660.1460.166,\;0.1460.166 , 0.146
$-5\bar{\sigma}_{{\rm H}}$ $0.723(4)$ 0.155, 0.1390.1550.1390.155,\;0.1390.155 , 0.139
Gauss + BB Numeric FC Cheat 0 ${\bf 0.688}(5)$ 0.148, 0.1300.1480.1300.148,\;0.1300.148 , 0.130
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.681}(5)$ 0.160, 0.1370.1600.1370.160,\;0.1370.160 , 0.137
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.685}(5)$ 0.149, 0.1290.1490.1290.149,\;0.1290.149 , 0.129
Gauss + BB Numeric Heuristic 0 ${\bf 0.675}(5)$ 0.145, 0.1440.1450.1440.145,\;0.1440.145 , 0.144
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.687}(5)$ 0.152, 0.1510.1520.1510.152,\;0.1510.152 , 0.151
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.688}(5)$ 0.153, 0.1520.1530.1520.153,\;0.1520.153 , 0.152
Table 10: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=20$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.030, $,\;0.030$ 0.030
$+5\sigma_{\infty}$ $0.683$ 0.030, $,\;0.030$ 0.030
$-5\sigma_{\infty}$ $0.683$ 0.030, $,\;0.030$ 0.030
Poisson Numeric Hessian 0 $0.518(5)$ 0.028, $,\;0.028$ 0.028
$+5\sigma_{\infty}$ $0.443(5)$ 0.028, 0.0270.0280.0270.028,\;0.0270.028 , 0.027
$-5\sigma_{\infty}$ $0.452(5)$ 0.028, $,\;0.028$ 0.028
Gauss Numeric Hessian 0 $0.518(5)$ 0.028, $,\;0.028$ 0.028
$+5\sigma_{\infty}$ $0.443(5)$ 0.028, 0.0270.0280.0270.028,\;0.0270.028 , 0.027
$-5\sigma_{\infty}$ $0.451(5)$ 0.028, $,\;0.028$ 0.028
Gauss Analytic Hessian 0 $0.518(5)$ 0.028, $,\;0.028$ 0.028
$+5\sigma_{\infty}$ $0.443(5)$ 0.028, 0.0270.0280.0270.028,\;0.0270.028 , 0.027
$-5\sigma_{\infty}$ $0.451(5)$ 0.028, $,\;0.028$ 0.028
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.690}(5)$ 0.039, $,\;0.039$ 0.039
$+5\sigma_{\infty}$ $0.604(5)$ 0.039, $,\;0.039$ 0.039
$-5\sigma_{\infty}$ $0.608(5)$ 0.039, $,\;0.039$ 0.039
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.690}(5)$ 0.039, $,\;0.039$ 0.039
$+5\sigma_{\infty}$ $0.604(5)$ 0.039, $,\;0.039$ 0.039
$-5\sigma_{\infty}$ $0.608(5)$ 0.039, $,\;0.039$ 0.039
Gauss + BB Numeric Hessian 0 $0.646(5)$ 0.042, $,\;0.042$ 0.042
$+5\sigma_{\infty}$ $0.642(5)$ 0.042, $,\;0.042$ 0.042
$-5\sigma_{\infty}$ $0.644(5)$ 0.042, $,\;0.042$ 0.042
Gauss + BB Numeric PLR 0 $0.647(5)$ 0.042, $,\;0.042$ 0.042
$+5\sigma_{\infty}$ $0.643(5)$ 0.043, $,\;0.043$ 0.043
$-5\sigma_{\infty}$ $0.646(5)$ 0.042, $,\;0.042$ 0.042
Gauss + BB Numeric sPFC 0 $0.518(5)$ 0.032, $,\;0.032$ 0.032
$+5\sigma_{\infty}$ $0.521(5)$ 0.032, $,\;0.032$ 0.032
$-5\sigma_{\infty}$ $0.520(5)$ 0.032, $,\;0.032$ 0.032
Gauss + BB Numeric PFC 0 $0.579(5)$ 0.039, 0.0360.0390.0360.039,\;0.0360.039 , 0.036
$+5\sigma_{\infty}$ $0.599(5)$ 0.040, 0.0360.0400.0360.040,\;0.0360.040 , 0.036
$-5\sigma_{\infty}$ $0.594(5)$ 0.040, 0.0360.0400.0360.040,\;0.0360.040 , 0.036
Gauss + BB Numeric Barlett 0 $0.571(5)$ 0.038, 0.0350.0380.0350.038,\;0.0350.038 , 0.035
$+5\sigma_{\infty}$ $0.582(5)$ 0.039, 0.0360.0390.0360.039,\;0.0360.039 , 0.036
$-5\sigma_{\infty}$ $0.576(5)$ 0.039, 0.0360.0390.0360.039,\;0.0360.039 , 0.036
Gauss + BB Numeric CH 0 ${\bf 0.693}(5)$ 0.046, 0.0430.0460.0430.046,\;0.0430.046 , 0.043
$+5\sigma_{\infty}$ $0.737(4)$ 0.048, 0.0440.0480.0440.048,\;0.0440.048 , 0.044
$-5\sigma_{\infty}$ $0.729(4)$ 0.048, 0.0440.0480.0440.048,\;0.0440.048 , 0.044
Gauss + BB Numeric FC Cheat 0 ${\bf 0.687}(5)$ 0.046, 0.0400.0460.0400.046,\;0.0400.046 , 0.040
$+5\sigma_{\infty}$ ${\bf 0.677}(5)$ 0.046, 0.0400.0460.0400.046,\;0.0400.046 , 0.040
$-5\sigma_{\infty}$ ${\bf 0.685}(5)$ 0.046, 0.0400.0460.0400.046,\;0.0400.046 , 0.040
Gauss + BB Numeric Ad hoc 0 ${\bf 0.690}(5)$ 0.046, $,\;0.046$ 0.046
$+5\sigma_{\infty}$ ${\bf 0.682}(5)$ 0.046, $,\;0.046$ 0.046
$-5\sigma_{\infty}$ ${\bf 0.681}(5)$ 0.046, $,\;0.046$ 0.046
Table 11: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{7}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Likelihood Minimim. CI method $\mu_{\rm t}$ Coverage $\hat{\sigma}$ (mean, median)
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.009, $,\;0.009$ 0.009
$+5\sigma_{\infty}$ $0.683$ 0.009, $,\;0.009$ 0.009
$-5\sigma_{\infty}$ $0.683$ 0.009, $,\;0.009$ 0.009
Poisson Numeric Hessian 0 $0.529(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
$+5\sigma_{\infty}$ $0.516(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
$-5\sigma_{\infty}$ $0.514(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
Gauss Numeric Hessian 0 $0.529(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
$+5\sigma_{\infty}$ $0.516(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
$-5\sigma_{\infty}$ $0.514(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
Gauss Analytic Hessian 0 $0.529(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
$+5\sigma_{\infty}$ $0.516(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
$-5\sigma_{\infty}$ $0.514(5)$ 0.009, 0.0100.0090.0100.009,\;0.0100.009 , 0.010
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.693}(5)$ 0.013, $,\;0.013$ 0.013
$+5\sigma_{\infty}$ ${\bf 0.678}(5)$ 0.013, $,\;0.013$ 0.013
$-5\sigma_{\infty}$ ${\bf 0.679}(5)$ 0.013, $,\;0.013$ 0.013
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.693}(5)$ 0.013, $,\;0.013$ 0.013
$+5\sigma_{\infty}$ ${\bf 0.678}(5)$ 0.013, $,\;0.013$ 0.013
$-5\sigma_{\infty}$ ${\bf 0.678}(5)$ 0.013, $,\;0.013$ 0.013
Gauss + BB Numeric Hessian 0 ${\bf 0.688}(5)$ 0.013, $,\;0.013$ 0.013
$+5\sigma_{\infty}$ ${\bf 0.675}(5)$ 0.013, $,\;0.013$ 0.013
$-5\sigma_{\infty}$ ${\bf 0.674}(5)$ 0.013, $,\;0.013$ 0.013
Gauss + BB Numeric PLR 0 ${\bf 0.688}(5)$ 0.013, $,\;0.013$ 0.013
$+5\sigma_{\infty}$ ${\bf 0.675}(5)$ 0.013, $,\;0.013$ 0.013
$-5\sigma_{\infty}$ ${\bf 0.674}(5)$ 0.013, $,\;0.013$ 0.013
Gauss + BB Numeric sPFC 0 $0.518(5)$ 0.009, $,\;0.009$ 0.009
$+5\sigma_{\infty}$ $0.522(5)$ 0.010, $,\;0.010$ 0.010
$-5\sigma_{\infty}$ $0.518(5)$ 0.010, $,\;0.010$ 0.010
Gauss + BB Numeric PFC 0 $0.590(5)$ 0.012, 0.0110.0120.0110.012,\;0.0110.012 , 0.011
$+5\sigma_{\infty}$ $0.591(5)$ 0.012, 0.0110.0120.0110.012,\;0.0110.012 , 0.011
$-5\sigma_{\infty}$ $0.602(5)$ 0.012, 0.0110.0120.0110.012,\;0.0110.012 , 0.011
Gauss + BB Numeric Barlett 0 $0.583(5)$ 0.011, $,\;0.011$ 0.011
$+5\sigma_{\infty}$ $0.582(5)$ 0.011, $,\;0.011$ 0.011
$-5\sigma_{\infty}$ $0.593(5)$ 0.011, $,\;0.011$ 0.011
Gauss + BB Numeric CH 0 ${\bf 0.686}(5)$ 0.014, 0.0130.0140.0130.014,\;0.0130.014 , 0.013
$+5\sigma_{\infty}$ ${\bf 0.694}(5)$ 0.014, 0.0130.0140.0130.014,\;0.0130.014 , 0.013
$-5\sigma_{\infty}$ ${\bf 0.701}(5)$ 0.014, 0.0130.0140.0130.014,\;0.0130.014 , 0.013
Gauss + BB Numeric FC Cheat 0 ${\bf 0.683}(5)$ 0.013, 0.0120.0130.0120.013,\;0.0120.013 , 0.012
$+5\sigma_{\infty}$ ${\bf 0.671}(5)$ 0.014, 0.0120.0140.0120.014,\;0.0120.014 , 0.012
$-5\sigma_{\infty}$ ${\bf 0.680}(5)$ 0.013, 0.0120.0130.0120.013,\;0.0120.013 , 0.012
Gauss + BB Numeric Ad hoc 0 ${\bf 0.692}(5)$ 0.013, 0.0140.0130.0140.013,\;0.0140.013 , 0.014
$+5\sigma_{\infty}$ ${\bf 0.677}(5)$ 0.013, 0.0140.0130.0140.013,\;0.0140.013 , 0.014
$-5\sigma_{\infty}$ ${\bf 0.679}(5)$ 0.013, 0.0140.0130.0140.013,\;0.0140.013 , 0.014
Table 12: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{8}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Acknowledgements.
C.A.A., L.B., and D.B. acknowledge financial support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement N. 101001205).